A Complete Bibliography of Philosophy of Science: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 November 2014 Version 1.01

Title word cross-reference

2 [596]. c [391]. log[P (h/eb)/P (h/b)] [463]. q [391].

-Numbers [391].

1 [188]. 1800 [203]. 1830 [34]. 1886 [57]. 1925 [87]. 1985 [91].

2 [188]. 20th [596].

Abduction [719]. Abductive [487]. Absence [270]. Absolute [213, 186, 113]. Abstraction [328]. Accardi [12]. Acceleration [409]. Acceptance [269]. Accessibility [325]. Accidental [357]. Accordance [6]. Account [412, 649]. Accounts [43, 667]. Accurate [374]. Achievements [180]. Action [187, 721, 262]. Actions [6]. Acts [48]. Acupuncture [419]. Adaptation [415, 471]. Adaptational [624]. Adaptationism [339].

Adaptations [49]. Adequacy [465]. Adolf [369, 404]. Adrift [256]. Adult [510]. Advancement [346]. Advancing [24]. Aesthetics [424]. Against [27, 496, 775, 360]. Age [259]. Agency [539, 679]. Aggregativity [568]. Agree [443]. Aharonov [571, 778, 650]. AI [7]. Aim [734]. Aims [180]. Alan [187]. Alberto [296]. Alexander [316, 473]. Alfred [30]. Algebra [604]. Algebraic [329]. Algebras [195]. Algorithms [464]. Alison [204, 512]. Allan [476, 206, 73]. Alleged [256]. Allen [404]. Alternative [634, 779]. Alternatives [714]. Althusser [273]. Altruism [268]. Amateur [57]. Ambiguous [368]. American [57]. Amsterdam [91]. Analog [294]. Analogical [550]. Analogies [757, 777]. Analysis [702, 266, 722, 746]. Analyst [96]. Analytic [267]. Anderson [229]. Andrew [679, 520]. Animal [368]. Animals [99]. Anomaly [158]. Anthropic [278]. Anthropology [674]. Anthropomorphic [351]. Anthropos [278]. Anti [632]. Anti-Realism [632]. Antiquarians [57]. Antiquity [276]. Antirealism [482, 224]. Antirealist [444]. Antireductionism [558, 459]. Anything [432]. Application [695]. Applications [332]. Applied [310]. Applying [720]. Approach [423, 44, 23, 534, 656, 83, 174, 180, 290, 76]. Appropriateness [291]. Approximate [375, 147]. Archaeological [205]. Archaeologists [57]. Archaeology [204, 487, 777, 204]. Architectonic [39]. Architecture [373, 438]. Argue [528]. Argument [148, 738, 52, 516, 179, 681, 146, 67, 775]. Arguments [774, 592, 68, 239, 223]. Aristotle [297, 73]. Ark [104]. Armstrong [212]. Arnold [654]. Aronson [182]. Arrow [493, 271]. Artificial [484, 556]. Ascribing [99]. Ascriptions [624]. Aspects [168, 281]. Assignment [82]. Association [449]. Astronomy [202, 224, 366]. Astrophysics [202]. Asymmetric [112]. Asymmetry [194, 251]. Attached [721]. Attitudes [656]. Attractors [485]. Attribution [99]. Autobiography [88]. Autonomy [63]. Availability [465]. Averted [447]. Away [429]. Axiology [708]. Axiom [12]. Axiomatization [254].

B [301, 56]. Bachelard [36]. Back [18, 41, 58, 77, 140, 93, 107, 127, 153, 169, 190, 214, 230, 246, 264, 285, 305, 320, 334, 352, 371, 394, 406, 416, 478, 489, 505, 420, 522, 525, 585, 598, 612, 625, 638, 657, 668, 685, 689, 740, 750, 771, 780, 593]. Backward [223]. Backward-Induction [223]. Bad [701]. Baker [656]. Balme [73]. Balzer [39]. Bang [272]. Barbara [389]. Barcan [386]. Barut [30]. Bas [275]. Based [174]. Bases [114]. Bayes [695, 739, 323, 588, 347]. Bayesian [347, 423, 693, 424, 193, 704, 707, 589]. Bayesianism [131, 238, 739, 520, 572, 521, 362]. Bayesians [443]. Be [432, 776, 443, 365]. Beamtimes [124]. Bechtel [54]. Becoming [515, 259]. Been [392]. Beginning [172, 66]. Behavior [410, 643, 641]. Behavioral [8, 116, 645]. Being [429, 259]. Belief [188, 773, 589]. Believing [492, 315]. Bell [628, 696, 145, 134]. Berkeley [89]. Bernstein [728]. Bertrand [288]. Best [705, 649, 284, 660, 466, 722]. Best-System [649, 722]. Better [577].

Betting [388]. Between [349, 619, 43, 413, 651, 90]. Beyond [699, 412, 739, 588]. Bias [456]. Bieder [57]. Big [272, 233]. Biochemistry [180, 746]. Biochemistry/Molecular [180]. Biological [415, 430, 475, 551, 340, 495]. Biologists [528]. Biology [339, 529, 530, 769, 477, 61, 555, 180, 497, 624, 562, 125, 473, 74, 497]. Biology/Cell [180]. Birthday [73]. Bishop [89]. Bloodsucking [351]. Bloor [226]. Blueprints [532]. Bob [653]. Bohm [571, 650, 744, 499, 778, 403, 311]. Bohr [37, 300]. Book [33, 318, 474, 762, 768, 168, 262, 273, 317, 92, 476, 37, 281, 152, 319, 682, 258, 770, 35, 167, 243, 228, 17, 277, 300, 610, 53, 500, 29, 72, 608, 763, 274, 636, 475, 761, 769, 74, 89, 103, 166, 87, 32, 75, 34, 226, 244, 208, 261, 477, 16, 202, 366, 245, 584, 36, 187, 206, 764, 607, 350, 654, 276, 122, 88, 284, 105, 501, 488, 405, 583, 188, 55, 203, 316, 382, 301, 369, 580, 656, 39, 304, 388, 348, 210, 346, 278, 503, 767, 211, 349, 391, 283, 592, 331, 165, 367, 387]. Book [212, 260, 106, 653, 766, 472, 91, 257, 681, 655, 123, 332, 139, 315, 38, 683, 204, 404, 282, 298, 279, 680, 302, 385, 259, 54, 213, 124, 296, 471, 651, 351, 497, 227, 383, 384, 205, 368, 679, 207, 609, 229, 765, 684, 209, 611, 652, 581, 504, 275, 386, 31, 579, 637, 90, 582, 30, 370, 76, 389, 393, 15, 299, 392, 297, 125, 502, 56, 73, 14, 303, 473, 280, 390, 347, 57]. Books [40, 126, 189, 263, 333]. Bootstrapping [65, 412, 377]. Border [38]. Born [637]. Boundaries [770]. Bourgeois [102]. Bowler [227, 227]. Bradshaw [32]. Brahe [202]. Brain [158, 552, 16, 501, 472]. Brainchildren [765]. Brandon [330, 471]. Breaking [401, 445]. Brian [188]. Britain [203]. Brown [208, 105, 387, 209, 15]. Bruce [205]. Bruna [245]. Bruno [187, 392]. Brussels [493]. Bub [763]. Buchwald [474]. Buckets [232]. Buckle [302]. Bust [593, 347]. Butts [587].

C [35, 17, 103, 764, 276, 39, 106, 765, 275]. Calculus [575, 203]. Campbell [699]. Can [432, 776, 558]. Canalization [422]. Capacities [618, 304]. Card [548]. Career [367]. Carl [752]. Carla [331]. Carnap [296]. Cartesian [365, 607]. Cartwright [651, 304]. Carving [222]. Case [691, 496, 540, 233, 116, 310, 458, 674]. Cat [651]. Category [715]. Cauchy [445]. Causal [375, 469, 700, 251, 26, 293, 138, 48, 345, 291, 280]. Causality [157, 381, 69, 1, 270, 314, 603, 249, 33, 764]. Causally [144]. Causation [171, 433, 706, 539, 163, 188, 716, 235, 270, 188, 636, 188]. Cause [623, 663, 728]. Causes [647]. Caveat [538]. Celebration [35]. Centenary [35]. Central [413, 27]. Century [680, 596, 752, 276]. Chain [297]. Challenge [339, 642, 619]. Challengeability [14]. Chalmers [754]. Chance [706, 649, 188, 384]. Chances [611, 280]. Change [396, 188, 290, 297, 303, 17]. Changing [700]. Chaos [219, 673, 630, 581, 161]. Character [491, 737, 292, 454]. Cheers [397]. Chemical [20]. Child [507, 512, 509, 511]. Children [455, 776]. Christa [87]. Christian [262]. Christopher [318]. Churchland [472, 501]. Circle [383]. Circles [401]. Civility [680]. Cladistic [309]. Clark [653]. Classical [391, 567, 729]. Classicism [438]. Classicism-Connectionism [438].

Classification [309]. Clearer [576]. Clerk [123]. Clinical [369]. Close [277]. Coffa [296]. Cognition [579]. Cognitive [742, 616, 776, 373, 313, 584, 276, 378, 557, 76, 389]. Coherence [184, 239]. Coherent [705]. Cold [277]. Cole [349, 350]. Collins [405]. Color [439, 106]. Colors [428, 439]. Colour [432]. Combinatorial [212]. Combining [634]. Comment [661, 365]. Commentary [643, 512]. Commentators [513, 662]. Comments [530, 739, 546, 498, 256]. Commitment [80]. Commitments [773]. Common [663, 728]. Communicative [262]. Comparative [419]. Competence [376]. Competition [447]. Complete [370]. Completeness [433]. Completing [43]. Complex [99]. Complexity [530, 555, 556, 746, 655]. Computable [161]. Computational [32, 472, 221]. Compute [558]. Concept [99, 340, 729]. Conception [2, 713, 135]. Concepts [339, 281, 172, 227, 495]. Conceptual [396, 7, 96, 20]. concerning [102, 135]. Conclusion [443]. Condition [758]. Conditionals [167, 681, 670]. Conditioning [601]. Conference [188]. Confirmation [152, 707, 673, 480, 116, 463, 758, 460, 347]. Confirmational [193]. Conflict [773]. Confusion [449]. Conjunction [159]. Connectionism [438]. Consciousness [262, 754, 614, 220, 379, 250]. Consensus [424]. Consequence [758]. Consequences [561]. Conservativeness [697]. Conserved [157, 381, 706]. Conspiracy [45]. Constraint [339]. Constraints [225]. Constructing [43, 654, 279]. Construction [348, 355]. Constructive [606]. Contemporary [538, 204]. Content [469, 228, 254, 50, 235, 327, 345, 236]. Content-Independent [327]. Context [469, 281]. Contexts [770, 551]. Contextualism [604]. Contingency [502]. Continua [766]. Continuing [635]. Contrastive [664, 138]. Control [516, 605]. Controversies [210]. Controversy [368]. Conventionality [414, 743, 312]. Converse [758]. Cooperative [717]. Copenhagen [502, 482]. Core [508]. Corpuscularian [89]. Correction [121]. Correlation [336]. Correlations [144]. Cosmological [278]. Cosmology [499, 732, 408, 172, 409, 66]. Cosmos [278]. Counterexample [623]. Counterfactual [270]. Counterfactuals [745, 447, 314, 260]. Couplings [461]. Covariant [329]. Creation [474, 272, 172]. Creative [32, 299]. Credence [188]. Credibility [688]. Crib [509]. Criteria [118]. Critical [732, 760, 759, 104, 746, 338, 135, 347, 204]. Criticisms [86]. Critique [2, 624]. Critiques [603]. Curtis [202]. Curve [423, 693, 628]. Cushing [502]. Cyborgs [390]. Cyclical [572].

D [243, 636, 517, 39, 304, 212]. Dangerous [678]. Daniel [317, 316, 765, 684]. Danilo [243]. Darden [303]. Darrigol [391]. Darwin [348, 678, 179, 68, 336, 301]. Darwinian [227, 464, 47]. Darwinism [437, 47]. Data [427, 401, 269]. Data-Technique [401]. Date [695]. David [92, 770, 167, 226, 206, 298, 73, 145]. David-Hillel [298]. Dawson [302]. Day [218, 611]. Dealing [755]. Dear [610]. Debate [438, 312, 282, 383]. Deborah [652]. Decade [160]. Decades [121, 103]. December [91].

Decision [188, 13, 48, 336, 6, 726, 682]. Decomposition [8]. Deduction [675]. Deductive [50]. Deductivism [70, 254, 139]. Deductivist [698]. Deep [290]. Defence [580]. Defending [719]. Defense [318, 573, 183, 96, 731]. Defensible [377]. Definability [173]. Defining [219]. Degrees [745]. Dembski [760]. Dennett [678, 765]. Dependence [171, 700]. Dependent [462]. Depragmatized [592]. Depth [291]. Descartes [350, 367, 736]. Description [110]. Design [415, 760, 746, 86]. Designing [765]. Detect [760]. Determinism [491, 400]. Development [281, 776, 203, 81]. Developmental [339, 643, 8, 641]. Developmentalist [642]. Diachronic [150]. Did [548, 101, 743]. Diderot [229]. Different [466]. Dilation [601]. Dim [397]. Dimensions [248]. Diminishing [316]. Direct [379]. Directedness [198]. Direction [527, 531]. Disappearance [22]. Discipline [610]. Disciplines [54]. Discourse [109]. Discover [552]. Discoverer [617]. Discovering [56]. Discovery [32, 477]. Discussion [635]. Disorder [393]. Dispositions [9, 633, 364]. Dispute [38]. Dissolution [738]. Disunity [770, 256, 393, 473]. Diverse [520, 521, 362]. Diversity [519, 644]. Divine [259]. Divisive [601]. DNP [623]. Do [688, 528, 589]. Does [425, 529, 28, 148, 186, 311]. Dogma [691]. Dogmas [60]. Donahue [366]. Donna [390]. Door [363]. Dream [370, 229]. Dreams [350]. Drift [448]. Dugald [37]. Duhem [60, 183, 315, 342, 589]. Duhemian [483]. Dupre [393, 1]. Dutch [592]. Dynamic [581, 239]. Dynamical [100]. Dynamics [279].

Early [57]. Earman [404, 213, 347]. East [419]. Ebbs [683]. Eckardt [389]. Ecological [734]. Ecology [8, 310]. Econometric [258]. Economic [316, 279, 684]. Economics [538, 91, 548, 207, 559, 317, 258, 207, 316]. Economist [559]. Edition [226]. Edwards [332]. Edwin [579]. Eells [69]. Effect [571, 650]. Effectiveness [327]. Effects [474, 678, 96]. Efficacy [517]. Ehrlich [766]. Einstein [72, 87, 111, 329, 270]. Either [491]. Electric [474]. Electricity [368]. Electromagnetic [123]. Elimination [242, 98]. Eliminative [200]. Eliminativism [11]. Eliminativist [446]. Elisabeth [152]. Ellery [69]. Elliot [497]. Elliott [475]. Embedding [85]. Embryo [558, 302]. Emerge [570]. Emergence [544, 717, 485, 557, 227, 383, 568]. Emergentism [643, 641]. Emerging [540]. Emotion [616]. Empirical [465, 363, 517, 341, 411, 619, 454]. Empiricism [606, 697, 363, 435, 117, 60, 101, 648, 752, 567]. Emptor [538]. Encoding [290]. Encounters [57]. End [703]. Endgame [109]. Energy [143, 124]. Engine [501]. Engineers [187]. England [680, 57]. Enlightenment [560]. Enough [647, 213]. Entail [186]. Entails [113]. Entering [563]. Environment [471]. Episode [596]. Epistemic [449]. Epistemological [555, 351]. Epistemology [193, 699, 243, 496, 266, 665, 3, 344, 289, 56, 731]. EPR [111, 144, 163, 604]. Epsilon [677]. Epsilon-Ergodicity [677]. Equal [28]. Equilibrium [640, 677, 729]. Equivalence [341]. Equivalent [270]. Eraser [467]. Ergodicity [677]. Ernst [769, 74]. Errol [278]. Error [541, 547, 589, 652]. Errors [234]. Erwin [605]. Essays [762, 475, 204, 404, 765, 316, 273, 228, 385, 386]. Essential [574]. Essentialism [136]. Establishing [534]. Ethnology [57]. Evaluation [518, 516, 605, 593, 378]. Evaluations [517]. Events [694]. Ever [365]. Every [611]. Everyone [405]. Evidence [322, 481, 529, 43, 413, 48, 129, 520, 146, 449, 521, 622, 747, 362]. Evidential [412, 519]. Evolution [196, 162, 631, 84, 439, 26, 348, 211, 21, 755]. Evolutionary [168, 491, 573, 152, 699, 475, 266, 737, 115, 292, 268, 624, 336, 176]. Exact [319]. Examination [347]. Excluded [447]. Exercise [483]. Experience [610]. Experientia [567]. Experiment [206, 540, 206]. Experimental [529, 401, 534, 748, 359, 555, 652]. Experimentation [530, 563, 566, 302]. Experiments [692, 774, 387, 86]. Expertise [536]. Explain [144, 234]. Explained [149]. Explaining [717, 656, 298, 76]. Explan [121]. Explanation [121, 375, 426, 705, 477, 623, 309, 577, 711, 356, 484, 138, 660, 378, 221, 327, 466, 603, 237, 674, 727, 10, 622, 747, 280, 569, 103, 103, 764, 284, 298]. Explanations [631, 664, 517, 543, 444, 721]. Explanatory [194, 109, 710, 175, 251, 380, 593, 725]. Explication [325]. Explorations [32]. Exploratory [563]. Exploring [564]. Exposing [442]. Extension [175]. Extensive [634]. Extent [725]. External [404]. Extraterrestrial [233].

F [477, 643, 88]. Faces [602, 728]. Factoring [593]. Facts [636, 723, 611]. Fail [552]. Failed [620]. Failure [337, 267]. Fair [755]. Fair-Dealing [755]. Fall [255]. Fälschung [72]. False [428]. Falsification [266]. Fascinating [611]. Fast [727]. Faye [300]. Feature [52]. Feminist [344, 56]. Fetzer [33]. Feyerabend [533, 545, 553, 567]. Field [694, 700, 105, 324, 488, 329, 134, 62, 635, 123]. Fields [563]. Fifth [12]. Finding [568]. Fine [104]. Firewalking [242]. Fisher [550]. Fitness [49, 495]. Fitting [423, 693]. Fleck [651]. Flight [629]. Fodor [228, 236]. Folgen [72]. Folk [373]. Follow [187]. Following [683]. Force [409, 130]. Foregone [443]. Forever [218]. Form [82]. Fortunes [75]. Foundational [438]. Foundations [167, 434, 34, 105, 411, 384, 393]. Four [121, 103]. Fraassen [102, 738, 468, 718, 451, 402, 275]. Framework [580]. Framing [561]. France [187]. Francois [88]. Frank [277, 331, 637]. Franklin [206, 240]. Frederick [661]. Frequency [174]. Freud [301, 148, 516, 370]. Friedman [319]. Frog [368]. Front [19, 42, 59, 79, 142, 94, 108, 128, 154, 170, 191, 216, 231, 247, 265, 287, 306, 321, 335, 354, 372, 395, 407, 418, 479, 490, 506, 421, 524, 526, 586, 599, 613, 627, 639, 658, 669, 687, 690, 741, 751, 772, 782]. Function [415, 655, 724]. Functional [309, 552, 721, 624, 237]. Functionalism [723]. Functionalist [64]. Functions [573, 198, 173, 225, 96]. Fundamental [619]. Further [601]. Fusion [277]. Future [8, 241, 95].

G [351, 205, 752]. Gabriel [89]. Galavotti [331]. Galison [770]. Galvani [368]. Galvanism [566]. Game [631, 336, 104]. Gary [32, 382, 683]. Gauge [445, 730]. Gauges [778]. Gavroglu [281]. Gazzaniga [584]. Geison [504]. Gene [534]. General [202, 425]. Generalizations [575, 132, 766]. Generalizing [268]. Genes [643, 641, 307]. Genetics [199, 645, 459, 303]. Genotype [534]. Genotype/Phenotype [534]. Geology [34]. Geometry [43, 713, 155, 31]. Gerald [404, 504]. Gereon [72]. German [315]. Germany [566]. Gestalt [182]. Getting [225]. Giere [76]. Giorgio [245]. Glance [559]. Goal [198]. Gödel [156]. Godfrey [655]. Godfrey-Smith [655]. Goldblatt [31]. Goldman [559]. Goldstick [119]. Golem [405]. Gooding [206]. Gopnik [512]. Got [589]. Gotthelf [73]. Goudaroulis [281]. Gravitational [440]. Gravity [271, 730]. Greenwood [518, 517, 183]. Grene [583]. Ground [359, 402]. Group [640, 514]. Growth [652, 343]. Grunbaum [404, 139, 605, 369]. GRW [701]. Guicciardini [203].

H [33, 167, 636, 366, 581, 596]. Habermas [262]. Hacking [359, 402, 582]. Hale [653]. Hall [649]. Hand [543, 245]. Handle [277]. Hanson [182]. Haraway [390]. Hard [357, 727]. Hardin [106]. Harding [56]. Harper [188]. Harre [182, 105]. Harris [278]. Harry [405]. Harvey [53, 105]. Hatfield [382, 365]. Hattiangadi [211]. Haugeland [762]. Hausman [317, 316]. Having [561, 762]. Healey [778, 650, 621]. Heat [207]. Hegemony [502]. Heidegger [63]. Heinrich [474]. Heisenberg [63]. Helen [244]. Helga [302]. Hellman [739]. Helmer [596]. Helmholtz [713]. Hempel [70, 596, 752]. Hempelian [758]. Hendry [123]. Herbert [32]. Hereditarian [227]. Heritability [249]. Heritage [300]. Heroic [459]. Hersh [767]. Hertz [474]. Heuristic [396]. Heuristics [568]. Hidden [491, 45, 134]. Hierarchy [308]. High [124]. Higher [84, 666]. Hilary [29]. Hill [318]. Hillel [298]. Hintikka [56]. Historians [57]. Historical [258, 358, 663, 555, 227, 560, 502, 73, 276]. History [258, 160, 208, 202, 654, 675, 391, 315, 204, 680, 596, 205, 458]. Hogarth [218]. Hole [452]. Holes [400, 195]. Holism [193]. Hollingdale [90]. Horwich [166, 519, 70]. Hot [277]. Hour [695]. Howson [240, 595]. Huggett [635]. Hugo [351]. Hull [92, 266]. Human [196, 8, 177]. Humphreys [280]. Hunger [565]. Husserl [83]. Hutchins [579]. Hydrogeology [561]. Hypotheses [24, 458]. Hypothesis [71]. Hypothetical [185]. Hypothetico [70, 254, 50]. Hypothetico-Deductive [50]. Hypothetico-Deductivism [70, 254].

Ian [582]. Idea [678]. Ideal [146, 326]. Ideas [258, 89]. Identifying [620]. Ideological [348]. If [167]. Igal [260]. Ihre [72]. II [72, 747]. Illusion [672, 346]. Image [698, 102]. Imagery [578, 226, 282]. Immune [234]. Imperatives [185]. Imperfect [326]. Implications [358, 344]. Imply [28]. Impossibility [374]. Impulse [328]. Incomplete [704]. Incompleteness [51, 122]. Inconsistency [606]. Inconsistent [408]. Independence [539, 293]. Independent [327]. Indeterminism [737, 445]. Indeterministic [491, 433]. Indexical [464, 722]. Indian [57]. Indices [197]. Indispensability [460]. Individual [110]. Individualism [573, 607, 11, 81, 237, 10, 779, 291]. Individuals [620]. Individuation [696]. Indivisible [641]. Induction [7, 575, 591, 233, 450, 192, 617, 223]. Inductive [575, 733, 681, 25]. Inequalities [134]. Inexact [317]. Infatuation [283]. Inference [705, 537, 681, 466, 112, 727, 760, 284]. Inferences [440]. Infinite [442, 503]. Inflationary [732]. Influence [301]. Information [78, 141, 215, 286, 353, 417, 523, 626, 686, 781, 665]. Informity [60]. Ingrao [245]. Inheritance [168]. Innate [376]. Innateness [422]. Inquiry [401, 238, 75, 244]. Instrumental [536, 457, 473]. Instrumentalism [199, 184]. Instrumentation [540]. Integrating [54]. Intellectual [87]. Intelligence [233]. Intelligent [746]. Intentional [721]. Intentionality [602]. Interdisciplinary [370]. Internal [404]. Interplay [540]. Interpolation [711]. Interpretation [744, 425, 427, 441, 468, 278, 62, 621]. Interpretations [434, 324]. Interpreting [763]. Interpretive [488, 635]. Intertheory [436]. Intervening [623]. Intervention [569]. Intimate [208]. Intimations [608]. Introduction [488, 635]. Intuitions [620]. Invariance [716, 445, 615, 569]. Investigations [566]. Invisible [245, 543]. Invisible-Hand [543]. Irrelevance [65]. Irreversibility [44, 135]. Irvine [188]. Irving [550]. Isaac [681]. Isomorphism [691]. Israel [245]. Issues [438, 384]. Itamar [165].

J [33, 770, 16, 348, 211, 472, 404, 227, 90, 14]. Jacob [88]. Jaegwon [385, 759]. James [208, 387, 209, 15, 502, 123]. Jan [300, 32]. Janis [404]. Jardine [75]. Jed [474]. Jeffrey [763, 155]. Jeffreys [575]. Jerry [228]. Johannes [366]. John [762, 261, 187, 350, 123, 315, 38, 404, 213, 351, 393, 145, 1, 517]. Joints [222]. Jonathan [368]. Jones [676]. Jordi [651]. Joseph [17, 39]. Journey [501]. Judgement [323]. Juhl [594]. Jungnickel [87]. Jurgen [262, 208]. Justice [631]. Justification [192].

K. [643]. Kant [296, 319, 257]. Kaplan [682]. Karen [302]. Karl [580]. Kasimba [302]. Kellert [581]. Kelly [594]. Kenneth [477]. Kepler [366]. Key [210]. Keynes [146]. Kill [101]. Kilmister [35]. Kim [759, 385]. Kincaid [180]. Kinds [437, 325]. Kitcher [103, 346, 257, 370]. Know [405]. Knowledge [508, 699, 22, 244, 654, 343, 168, 211, 279, 652, 226]. Knows [695]. Kosso [500]. Kuhn [597, 101]. Kuhse [302]. Kurt [156]. Kvart [260].

L [92, 32, 188, 106, 504]. Laboratory [387]. Landé [30]. Langley [32]. Language [492, 715, 211]. Languages [361]. Large [174]. Larry [210, 611]. Latour [187, 392]. Laudan [34, 708, 210, 595, 611]. Lavine [503]. Law [323, 575, 649, 174, 449, 187]. Law-Like [575]. Lawhood [745]. Lawlessness [562]. Lawrence [384]. Laws [529, 532, 430, 633, 745, 549, 325, 722, 275].

Leaps [447]. Learning [590, 547]. Leeds [621]. Legacy [243, 91, 300]. Legend [346]. Leibniz [195]. Lenhardt [262]. Lepenies [90]. Leplin [137, 136]. Leslie [261]. Level [438]. Levels [330]. Levi [13, 681]. Levine [57]. Lewis [145, 649, 722]. Lewisian [270]. Life [484]. Lifetimes [124]. Light [703, 207]. Like [575, 528]. Likelihood [692]. Limitations [139]. Limiting [174]. Limits [735]. Lindley [303]. Linear [155]. Lines [447]. Linguistics [663]. Linking [742, 251]. Linnaean [308]. Lipton [284]. Literature [90]. Little [684]. Living [769, 73]. Lloyd [152]. Local [329]. Locality [608, 45, 270]. Logic [591, 733, 38, 310, 336, 165]. Logical [434, 82, 118, 101, 648, 383, 752]. Lola [651]. Long [68]. Longino [244, 731]. Look [732, 400]. Lorentz [716, 329]. Lorentz-Covariant [329]. Louis [273, 504]. Low [281]. Loyalty [536]. Lucille [301]. Lüders [110]. Lyell [47]. Lynne [656]. Lyon [315].

M [317, 258, 32, 501, 580, 212, 351, 609, 73]. Mach [72, 72]. Machinery [442]. Macnamara [38]. Macroscopic [483, 749]. Maddy [768, 274]. Made [705, 307, 495]. Magic [367]. Magnitudes [600]. Maher [481, 494, 594, 131, 388]. Major [267]. Make [629, 148]. Making [547, 349]. Malament [218, 743]. Mandelbaum [368]. Mangle [679]. Manufactures [674]. Marcello [368]. March [91]. Marcus [386]. Marjorie [583]. Mark [682]. Markov [112]. Martin [315]. Marxism [64]. Mary [36]. Massey [404]. Mastery [87]. Material [744]. Materialism [397, 318]. Maternal [678]. Mathematical [610, 316, 217, 376]. Mathematics [87, 715, 767, 720, 460, 768, 274, 331]. Matter [18, 41, 58, 77, 19, 42, 59, 79, 140, 93, 107, 127, 142, 94, 108, 128, 153, 169, 190, 214, 154, 170, 191, 216, 230, 246, 264, 285, 231, 247, 265, 287, 305, 320, 334, 352, 306, 321, 335, 354, 371, 394, 406, 416, 372, 395, 407, 418, 478, 489, 505, 420, 522, 525, 479, 490, 506, 421, 524, 526, 585, 598, 612, 625, 586, 599, 613, 627, 638, 657, 668, 685, 689, 639, 658, 669, 687, 690, 740, 750, 771, 780, 741, 751, 772, 782, 173]. Maudlin [608, 757]. Maximization [6]. Maxwell [123]. Mayo [652, 86]. Mayr [769, 74]. McCormmach [87]. McKie [243]. Meaning [646, 348, 137, 723]. Meant [365]. Measure [707, 463]. Measured [727]. Measurement [634, 363, 403, 718, 451, 311, 667, 621, 304]. Measurements [374, 332]. Measures [707]. Mechanics [465, 44, 640, 696, 754, 434, 414, 82, 51, 468, 255, 677, 502, 729, 122, 384]. Mechanism [26]. Medical [280]. Medicine [477]. Meets [145, 336]. Megill [476]. Mellor [636]. Membership [574]. Memory [582]. Mendeleev [131]. Mendelian [227, 303]. Mental [99, 158, 427, 578, 46, 98]. Mercenary [159]. Mere [272]. Merrill [56]. Merwe [30]. Mesosomes [748]. Metaphoric [379]. Metaphysical [608, 237, 393]. Metaphysics [762, 56]. Method [50, 256, 253, 137, 564]. Methodological [99, 2, 11, 185, 454, 10, 281]. Methodologist [148]. Methodology [56, 316]. Methods [619, 466, 178]. Meyering [276]. Mice [102]. Michael [319, 584, 122, 282, 125, 548]. Microphysics [554]. Microscope [688]. Middle [359]. Mighty [87]. Mill [545, 466]. Mind [158, 761, 759, 175, 387, 655, 552, 385, 27, 343, 759, 762, 607, 656, 370]. Mind-Brain [158]. Mind/Brain [175]. Minds [607, 765]. Mineralogy [34]. Mirowski [207]. Mirrors [527]. Misrepresentation [46]. Mistake [239]. Mittelstrass [208]. Modal [434, 441, 468, 621]. Modalities [386]. Model [23, 738, 623, 645]. Model-Theoretic [23, 738]. Modeling [430]. Models [532, 427, 542, 673, 550, 551, 684]. Modern [258, 17, 608, 227, 267, 14, 392]. Moked [89]. Molecular [61, 307]. Monetary [550]. Moon [672]. Moral [262]. Morgan [258]. Morphological [348]. Motion [367, 736]. Motor [578]. Moulines [39]. Multiple [742, 777, 775, 582]. Multiplicity [342]. Murdoch [37]. Myth [580, 564, 227].

N [211, 315, 471, 76]. Nancy [304, 651]. Narlikar [272]. Narrow [573]. Natural [415, 162, 254, 437, 206, 382, 387, 259, 556, 470]. Naturalism [5, 4, 383, 768]. Naturalist [2]. Naturalistic [496, 708, 269]. Naturalized [252, 713, 457, 178]. Nature [87, 80, 748, 222, 349, 207, 73, 722, 500, 655, 390, 304]. Necessity [547]. Need [425]. Neglect [714]. Neil [91]. Neo [60]. Neo-Empiricism [60]. Networks [533]. Neural [742, 533]. Neurath [651, 383, 269, 243]. Neurobiology [120]. Neuroimages [552]. Neuropsychology [220]. Neurosciences [584]. Never [392]. Newton [202]. Newtonian [203, 408, 409]. Nicholas [168, 75, 404]. Nicholsen [262]. Niels [37, 300]. Nietzsche [63]. Nineteenth [276]. No [695, 491, 443, 411, 721, 104]. No-Signalling [411]. NOA [104]. Noccolo [203]. Nomic [171]. Non [608, 664, 397, 227, 743]. Non-Conventionality [743]. Non-Darwinian [227]. Non-Locality [608]. Non-Probabilistic [664]. Non-Reductive [397]. Nonhuman [99]. Nonlocality [122, 571, 604, 716]. Nonmonotonic [681]. Nonpragmatic [671]. Nonseparability [696, 515, 630]. Normality [340]. Normative [5, 4, 382]. Norms [80]. Norton [609]. Note [716, 468, 467]. Notes [331]. Notice [760, 759, 104]. Notion [115]. Notturno [580]. Novel [129]. Novelty [714, 458]. Numbers [600, 391, 367, 233, 766, 174]. Nutini [351].

O [30, 14, 596]. Obituary [587]. Objections [236]. Objective [480]. Objectivity [401, 244, 346, 379, 476, 36]. Objects [694, 744, 483]. Observables [604]. Observation [703, 120, 628, 60, 164]. Observations [196]. Off [330, 151, 470, 700]. Ohm [87]. Old [622, 747]. Olivier [391]. Olympian [350]. Omniscient [448]. One [695, 26, 463, 47, 641, 459]. O’Neill [119]. Ontic [426]. Ontology [694, 483, 55]. Openness [241, 95]. Opens [363]. Operationism [399]. Operations [634]. Operators [51, 155]. Opinion [155]. Oppenheim [596]. Oracles [424]. Order [666, 581]. Organisms [49, 620, 645]. Organization [699]. Origin [179]. Origination [272]. Origins [358, 399, 379]. Orthogonality [31]. Osiris [654]. Other [273, 228]. Otto [243, 651]. Outbreaks [562]. Overcoming [383].

P [167, 257]. Paleobiology [710]. Paper [661, 660, 659]. Papers [756, 91, 30]. Paradigm [266]. Paradox [435, 733, 288, 146, 223, 728]. Part [202]. Particles [89]. Parts [552, 55]. Pascal [302]. Passage [28]. Pasteur [504]. Pasteurization [187]. Pat [32]. Patricia [472, 370]. Patrick [388]. Patterns [540, 566]. Paul [166, 501, 488, 280]. Peacocke [259]. Brain [175]. Cell [180]. Miller [52]. Molecular [180]. Phenotype [534]. Penelope [768, 274]. Pera [368]. Perception [691, 276, 379]. Perceptual [616, 290]. Personality [582]. Perspective [358]. Perspectives [258, 56]. ’Pessimistic [450]. PET [564]. Peter [770, 610, 500, 284, 55, 653, 655, 302, 227]. Phenomena [533, 102, 440, 717]. Phenomenal [753]. Phenomenological [619]. Phenomenology [156]. Phenomenotechnique [358]. Phenotypic [21]. Philip [103, 346, 766, 207, 88]. Philippa [57]. Philosophers [256, 106]. Philosophical [243, 501, 583, 278, 211, 385, 384, 560, 386, 338, 299, 73, 105, 404]. Philosophy [419, 273, 629, 37, 358, 295, 23, 252, 74, 208, 538, 122, 590, 316, 210, 546, 283, 331, 548, 83, 315, 204, 404, 344, 596, 651, 4, 497, 133, 752, 267, 457, 562, 178, 56, 682, 475, 89, 273, 16, 497, 125]. Phylogenetic [644]. Physical [9, 43, 759, 607, 551, 364, 495, 280]. Physicalism [735]. Physicalist [558]. Physicist [315]. Physics [281, 87, 440, 217, 207, 63, 297, 37, 608, 124, 299, 384]. Pickering [679]. Pieces [552]. Pierre [315, 315]. Pinch [206, 405]. Pinker [761]. Pinsky [204]. Pitowsky [218, 165]. Pitt [17]. Place [413, 38]. Placebo [516, 605]. Planetary [202]. Plantinga [496]. Plasticity [290]. Play [509, 511]. Playing [548]. Plumpton [331]. Pluralism [200, 632, 710, 545, 648, 360]. Plurality [707, 178]. Point [475]. Points [419, 537]. Polanyi [548]. Policy [561]. Politics [316, 204, 133, 565, 651]. Polymath [35]. Popper [580, 575, 52, 225, 146]. Popper/Miller [52]. Popperian [91]. Positing [27]. Positivism [383]. Possibility [212]. Possible [441, 211]. Postmodern [133]. Postulate [143]. Power [593, 770]. Powers [469, 345]. Practical [656]. Practice [453, 679, 456]. Pragmatic [549, 269]. Precision [609]. Prediction [240]. Predictivism [494]. Preface [431, 535, 712]. Preference [670]. Preferences [572]. Preparation [718, 451]. Presented [91, 73]. Prigogine [100, 135]. Principle [692, 663, 28, 728, 278]. Priori [365]. Private [504]. Probabilism [671]. Probabilistic [181, 69, 1, 664, 25, 270]. Probabilities [541]. Probability [100, 575, 82, 188, 173, 225, 331, 165, 174, 12, 33]. Probably [232]. Problem [423, 693, 194, 531, 707, 118, 445, 403, 134, 648, 452, 338, 311, 589]. Problems [99, 698, 308, 404, 236]. Proceedings [188]. Process [157, 776, 437, 641, 112, 92, 32]. Processes [99, 749]. Products [438]. Professional [57]. Program [195]. Progress [629, 17, 115, 727]. Prohibitions [270]. Projection [143]. Prolegomenon [122, 62]. Proliferation [533]. Promise [197]. Proof [491]. Proofs [411]. Proper [724, 62]. Properties [9, 570, 364]. Proposal [616]. Proposals [100]. Prospects [397]. Protective [363]. Protocol [383, 269]. Prove [743].

Pseudoscience [648]. Psychoanalysis [67, 369]. Psychological [132]. Psychology [573, 704, 373, 607, 727, 291, 257, 38]. Psychophysical [235]. Psychotherapeutic [517]. Psychotherapy [518, 516, 605]. Publicity [614]. Pure [709, 720]. Putnam [653, 29, 738].

Q [167]. Qualitative [116, 758]. Qualms [64]. Quantifiers [646]. Quantities [381, 600]. Quantity [157, 706]. Quantum [465, 694, 696, 754, 499, 700, 109, 363, 434, 763, 414, 82, 122, 757, 51, 105, 324, 488, 493, 411, 630, 468, 271, 391, 165, 329, 134, 467, 63, 255, 66, 62, 635, 12, 608, 165, 502]. Quark [458]. Quasi [697]. Quasi-Truth [697]. Questions [620, 561]. Quine [342, 60, 183].

R [208, 350, 105, 580, 315, 90, 709]. R-Selection [709]. Race [277]. Rachel [34]. Railton [623]. Raising [706]. Ramsey [331, 681, 646]. Random [448]. Randomization [86]. Randomness [100]. Rational [209]. Rationality [118, 150, 457, 177, 15, 580]. Ravens [733]. Reading [500, 653]. Real [766]. Realism [426, 363, 117, 632, 122, 97, 201, 553, 450, 359, 360, 727, 130, 147, 683, 274]. Realist [328]. Reality [379, 402, 56, 29]. Realizability [742, 775]. Really [709, 767, 408, 466]. Reals [766]. Reason [323, 501, 6]. Reasoning [698, 748, 443, 487, 177, 167, 681]. Reasons [337]. Rebel [637]. Rebellion [350]. Received [40, 126, 189, 263, 333]. Reciprocity [413]. Recognize [47]. Reconciling [616]. Reconsidering [729]. Reconstruction [348]. Records [465]. Recovering [446]. Recurrent [533]. Recursive [173]. Recycling [536]. Redei [700]. Redhead [122]. Redshift [540]. Reduced [432]. Reduction [242, 199, 175, 355, 98]. Reductionism [158, 356, 397, 775]. Reductionist [180]. Reductive [397, 568]. Redundant [746]. Reference [629, 615, 186, 113]. Reflections [211, 560, 299, 628]. Reflexive [243]. Refutation [567]. Refuted [53]. Regained [223]. Reinterpretation [722]. Reinterpreting [227]. Reinvention [390]. Rejoinder [693, 69, 645, 621]. Relation [43, 208, 436, 143]. Relational [213, 736]. Relationist [85]. Relationships [534]. Relative [409]. Relativism [210, 53]. Relativistic [700, 414, 666, 134]. Relativitätstheorie [72]. Relativity [608, 425, 28, 515, 241, 743, 95]. Relevance [412, 380]. Relevant [714]. Reliability [692, 217, 684]. Reliable [465, 238]. Reliance [159]. Remarks [555, 135]. Renaissance [202]. Rene [202, 350, 367]. Renormalization [640, 442]. Renormalizing [3]. Repeated [621]. Reply [481, 363, 381, 1, 518, 102, 513, 605, 517, 757, 675, 183, 520, 137, 603, 676, 662, 86]. Representation [464, 542, 326, 29]. Representational [616]. Rescher [168, 404]. Resolution [288]. Response [595, 180]. Responses [594]. Responsibility [546]. Rest [455]. Results [601]. Rethinking [476]. Retreat [553]. Return [514]. Returns [316]. Reuben [767]. Rev. [695]. Revealed [670]. Review [33, 318, 474, 762, 768, 168, 262, 273, 317, 92, 476, 37, 281, 152, 319, 682, 258, 770, 35, 167, 243, 228, 17, 277, 300, 610, 53, 500, 29, 72, 608, 763, 274, 636, 475, 761, 769, 74, 89, 166, 87, 32, 75, 34, 226, 244, 208, 261, 477, 16, 202, 366, 245, 584, 36, 187, 764, 607, 350, 654, 276, 122, 88, 284, 105, 501, 488, 405, 583, 55, 203, 382, 301, 369, 580, 656, 39, 304, 388, 348, 210, 346, 278, 503, 767, 211, 349, 391, 283, 331, 165, 367, 387, 212, 260, 106, 653]. Review [766, 472, 91, 257, 681, 655, 123, 332, 139, 38, 683, 204, 404, 282, 298, 279, 680, 302, 385, 259, 54, 213, 124, 296, 471, 651, 351, 497, 383, 384, 205, 368, 679, 207, 609, 229, 765, 684, 209, 611, 652, 581, 504, 275, 386, 338, 498, 31, 579, 637, 90, 582, 30, 370, 76, 635, 389, 393, 15, 299, 392, 297, 125, 502, 56, 73, 14, 303, 473, 280, 390, 347, 57]. Reviews [103, 206, 188, 316, 315, 227]. Revisited [742, 495]. Revolution [227, 20, 610]. Revolutions [117]. Rewriting [582]. Rhetoric [482]. Richards [348]. Right [206, 225]. Rings [195]. Rise [202, 276, 255, 90]. Risks [611]. Ritvo [301]. Robert [607, 348, 387, 209, 31, 15, 57, 587]. Roberto [299]. Roberts [351]. Robustness [163]. Role [546, 4, 752]. Rom [105]. Ronald [76]. Room [491]. Roots [276]. Rosenberg [316, 498, 473]. Roy [279]. Ruben [298]. Rudder [656]. Ruetsche [718]. Rule [110, 683]. Rule-Following [683]. Rules [185, 454]. Rural [351]. Ruse [125]. Russell [87, 71]. Ruth [386].

S [318, 597, 258, 584, 399, 367, 472]. S. [399]. Sake [681]. Salmon [103, 157, 381, 103, 160, 764, 380, 139]. Same [776]. Sandra [56]. Sanford [167]. Saving [102]. Say [701]. Schaffer [206]. Schaffner [477, 643, 624]. School [100, 493]. Schrödinger [35]. Schulte [594]. Science [419, 317, 647, 629, 508, 358, 295, 770, 23, 2, 252, 769, 509, 102, 34, 628, 614, 36, 538, 276, 753, 590, 61, 316, 580, 444, 346, 546, 349, 548, 378, 83, 680, 344, 365, 651, 133, 227, 752, 453, 104, 560, 457, 504, 90, 178, 370, 76, 389, 565, 149, 459, 460, 303, 57, 17, 208, 654, 405, 39, 210, 283, 315, 679, 393, 56, 14, 473, 92, 244, 187, 210]. Sciences [607, 175, 116, 327, 582, 319, 206, 301, 387, 280]. Scientific [121, 474, 342, 426, 328, 698, 610, 776, 117, 103, 661, 244, 313, 537, 756, 238, 484, 660, 367, 253, 447, 450, 259, 54, 596, 456, 177, 30, 659, 147, 103, 32, 15]. Scientism [283]. Scientist [510, 507, 512]. Scientists [273, 187]. Screening [330, 700, 151, 470]. Screening-Off [330, 151, 470]. Search [176]. Seat [501]. Second [226]. See [688]. Sejnowski [472]. Selected [96, 385, 30]. Selection [248, 162, 330, 266, 709, 26, 724, 21, 725, 151, 338, 470, 114, 130]. Selector [688]. Self [374]. Self-Measurements [374]. Semantic [83, 81, 296]. Semi [619]. Semi-Empirical [619]. Sensations [318]. Sensitivity [707]. Sentence [383]. Sentences [646]. Separability [773]. Separate [317]. Seventeenth [680]. Seventeenth-Century [680]. Seventieth [73]. Several [443]. Severe [129]. Severity [714, 458]. Sex [755]. Shaky [104]. Shapere [164]. Shapin [680]. Sharon [124]. Shaughan [503]. Shea [367]. Sheridan [187]. Shierry [262]. Shifts [155]. Short [68]. Should [144, 405]. Siegel [53]. Signalling [411]. Signals [45]. Signs [713]. Simians [390]. Similarity [436]. Simon [32, 206]. Simons [55]. Simple [162, 24, 644]. Simulation [578, 294]. Simultaneity [414, 743, 312]. Singer [302]. Singular [702]. Singularity [618]. Skepticism [575]. Sklar [384]. Skyrms [188, 188]. Sloth [118]. Smith [655]. Sneed [39]. Sober [330, 475, 497]. Social [688, 226, 244, 546, 116, 680, 207, 723, 280, 731, 209]. Socializing [252]. Society [187, 343, 227, 349]. Socio [204]. Socio-Politics [204]. Sociological [15]. Sociology [90]. Soft [357]. Sola [567]. Solar [540]. Solution [452]. Solve [311]. Solves [403]. Some [308, 620, 666, 210, 555, 560, 498, 135]. Sophisticated [148]. Sorell [283]. Soul [501, 582]. Source [737]. Space [43, 232, 213, 186, 113, 461]. Space-Time [43, 213, 461]. Spaces [7]. Spacetime [696, 232, 736, 31]. Spacetimes [702, 218]. Spatial [713]. Special [743]. Species [84, 632, 620, 486, 360]. Spectrometer [289]. Speech [691]. Spirit [482, 196, 752]. Split [438]. Split-Level [438]. Spontaneous [273]. Stability [745, 6]. Stabilizing [279]. Standards [2]. State [374, 110]. Statements [269]. States [742]. Station [296]. Statistical [44, 640, 664, 737, 292, 384, 727, 677, 729]. Statistics [188, 547, 589]. Statue [88]. Steel [675]. Stephen [349, 302, 581]. Steven [761, 680]. Stevens [399]. Stochastic [329, 270]. Strange [485, 461]. Strategies [303]. Strategy [176]. Stretched [447]. Strings [721]. Stronger [322]. Structural [325]. Structuralism [437]. Structure [152, 43, 661, 313, 659, 20, 734, 461]. Structures [756]. Studies [208, 206, 540, 310, 73]. Study [691, 748, 55, 351, 674, 297]. Stump [770]. Subject [700]. Subjective [480]. Subjectivity [379]. Subjunctive [670]. Substance [232]. Substantivalism [195]. Success [444, 677]. Sufficiency [181]. Sulloway [637]. Superluminal [45, 270]. Supernaturalism [351]. Superselection [667]. Supertasks [218, 666]. Supervenience [544, 221, 235, 495, 385]. Suppe [661]. Support [25]. Suspicious [52]. Swimming [481]. Switching [182]. Symmetrized [255]. Symmetry [445, 554, 275]. Symposium [91]. Symptomatic [48]. Syntactic [452]. Synthesis [176]. System [649, 234, 722]. Systematic [714]. Systems [341, 27, 644, 581].

T [502]. Tails [701]. Take [611]. Takes [429]. Tale [301, 459]. Tales [509]. Tally [148, 516, 67]. Targeting [534]. Taton [202]. Taxa [84]. Technique [401]. Technology [550]. Teller [488]. Tells [220]. Temperature [281]. Tempered [130]. Temporal [135]. Temporally [112]. Tensor [438]. Term [722]. Terms [462]. Terrence [472]. Terrible [759]. Test [253, 681]. Testability [132]. Testament [446, 583]. Testing [458]. Tests [342, 129]. Thackray [654]. Thau [649]. Their [304, 173]. Theo [276]. Theology [259]. Theorem [695, 696]. Theoretic [23, 631, 738]. Theoretical [99, 430, 87, 517, 80, 291]. Theories [100, 196, 616, 428, 254, 313, 337, 766, 332, 85, 213, 455, 388]. Theory [168, 694, 162, 491, 499, 43, 181, 682, 700, 228, 109, 157, 706, 266, 737, 60, 665, 276, 540, 292, 324, 713, 356, 493, 590, 369, 413, 715, 593, 13, 48, 212, 403, 260, 447, 123, 221, 329, 462, 134, 186, 743, 236, 336, 113, 311, 62, 159, 269, 12, 734, 730, 197, 241, 95, 152, 105, 488, 348, 391, 347, 635, 303]. Theory-Conjunction [159]. Theory-Dependent [462]. Theory-Informity [60]. There [443, 452, 47]. Thesis [60, 183]. Thing [759]. Things [393, 73]. Third [160]. Thomas [651, 383, 597]. Thought [762, 774, 361, 387, 205]. Thoughts [494]. Tiles [36]. Tim [608]. Time [527, 499, 531, 43, 143, 271, 554, 213, 186, 679, 255, 113, 461, 493, 213]. Time-Energy [143]. Time-Symmetrized [255]. Times [28]. Tlaxcala [351]. Today [125]. Tom [283]. Too [277]. Top [458]. Torch [87]. Torretti [299]. Tort [449]. Toxic [449]. Tradition [296]. Traditions [204]. Transcendental [257]. Transform [533]. Transformations [110]. Translation [375]. Traweek [124]. Treatments [516, 605]. Trevor [206, 405]. Triangulation [138]. Trigger [205]. Trouble [667]. True [428, 463]. Truer [119]. Trump [548]. Truth [697, 50, 680, 402, 147, 166]. Truthlikeness [375, 615]. Tunneling [703]. Turn [15]. Twentieth [752]. Twentieth-Century [752]. Twist [312]. Two [339, 602, 64, 537, 60, 301, 397, 603, 562]. Tycho [202]. Tye [282]. Type [318]. Types [49].

Uebel [651, 383]. Ulises [39]. Unanimity [181]. Unbounded [51]. Uncertainty [143]. Underdetermination [703, 341]. Understanding [647, 109, 456, 299, 503]. Unification [194, 356, 675, 725]. Uniformities [357]. Uniqueness [85]. Unit [26]. Units [410, 84, 21, 151, 338, 726, 114]. Unity [61]. Universal [575]. Universality [640, 618]. Universe [272, 66]. Universes [261]. Unorthodox [695]. Unpredictable [581]. Untenable [359]. Updated [398]. Urbach [86]. Use [665, 714, 453]. Use-Novelty [714]. Useful [168]. Uses [206, 563]. Utility [332].

Valerie [204]. Validation [67, 369]. Value [773, 48, 521]. Values [244, 609]. Variables [491, 134]. Variation [566]. Various [756]. Verification [566]. versus [213]. Vicarious [688]. Victorian [57]. Vienna [296, 383]. View [475, 537]. Viewpoint [448]. Views [624]. VII [297]. Vindication [671]. Virtual [402]. Virtue [547]. Vision [439, 221, 576]. Visual [487]. Visualizations [453]. Vol [87, 202]. Volta [368]. Volume [78, 141, 215, 286, 353, 417, 523, 626, 686, 781, 654, 188]. vs [402].

W [35, 139]. Wake [581]. Ward [332]. Wardy [297]. Warfare [674]. Was [272, 365]. Waste [759]. Water [232]. Waves [232, 474]. Way [610, 589, 47]. Wayne [519, 520]. Weak [174]. Weber [262]. Weight [146]. Weingard [635]. Weintraub [279]. Went [628]. Wesley [103, 764, 157]. West [419]. Western [674]. Wild [579]. Wilda [229]. Will [443]. William [366, 188, 367, 54, 760]. Wilson [607, 202]. Winnie [28]. Wisdom [398, 14]. Wise [609]. Witchcraft [351]. within [88, 383, 267]. without [439, 346, 314]. Wolf [90]. Wolfgang [39]. Wolter [72]. Women [390]. Work [548, 315, 338]. Works [640, 761]. World [429, 763, 341, 124, 572, 769, 759, 213]. Worlds [441, 550, 404]. Wormwholes [643]. Wrong [160, 266, 628, 543, 206]. Wylie [204].

Yang [778]. Years [57]. Young [16]. Youthful [350].

Z [474, 16]. Zolo [243]. Zytkow [32].

References

Eells:1987:PCR

[1] Ellery Eells. Probabilistic causality: Reply to John Dupré. Philosophy of Science, 54(1):105–114, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187477. See [69].

Doppelt:1990:NCM

[2] Gerald Doppelt. The naturalist conception of methodological standards in science: A critique. Philosophy of Science, 57(1):1–19, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187617.

Leplin:1990:RE

[3] Jarrett Leplin. Renormalizing epistemology. Philosophy of Science, 57(1): 20–33, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187618.

Rosenberg:1990:NNR

[4] Alexander Rosenberg. Normative naturalism and the role of philosophy. Philosophy of Science, 57(1):34–43, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187619.

Laudan:1990:NN

[5] Larry Laudan. Normative naturalism. Philosophy of Science, 57(1):44– 59, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187620.

Sobel:1990:MSD

[6] Jordan Howard Sobel. Maximization, stability of decision, and actions in accordance with reason. Philosophy of Science, 57(1):60–77, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187621.

Gardenfors:1990:ICS

[7] Peter Gärdenfors. Induction, conceptual spaces and AI. Philosophy of Science, 57(1):78–95, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187622.

Kitcher:1990:DDF

[8] Philip Kitcher. Developmental decomposition and the future of human behavioral ecology. Philosophy of Science, 57(1):96–117, March 1990. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187623.

Averill:1990:PPD

[9] Edward Wilson Averill. Are physical properties dispositions? Philosophy of Science, 57(1):118–132, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187624.

Tuomela:1990:MIE

[10] Raimo Tuomela. Methodological individualism and explanation. Philosophy of Science, 57(1):133–140, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187625.

Kincaid:1990:EMI

[11] Harold Kincaid. Eliminativism and methodological individualism. Philosophy of Science, 57(1):141–148, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187626.

VanDenBerg:1990:AQT

[12] Hans Van Den Berg, Dick Hoekzema, and Hans Radder. Accardi on quantum theory and the “fifth axiom” of probability. Philosophy of Science, 57(1):149–157, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187627.

Leeds:1990:LDT

[13] Stephen Leeds. Levi’s decision theory. Philosophy of Science, 57(1):158– 168, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187628.

Wilson:1990:BRB

[14] Fred Wilson. Book review: Challengeability in Modern Science by J. O. Wisdom. Philosophy of Science, 57(1):169–170, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187629.

Tibbetts:1990:BRB

[15] Paul Tibbetts. Book review: Scientific Rationality: The Sociological Turn by James Robert Brown. Philosophy of Science, 57(1):170–172, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187630.

Gillett:1990:BRB

[16] Grant Gillett. Book review: Philosophy and the Brain by J. Z. Young. Philosophy of Science, 57(1):172–173, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187631.

Cushing:1990:BRB

[17] James T. Cushing. Book review: Change and Progress in Modern Science by Joseph C. Pitt. Philosophy of Science, 57(1):173–176, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187632.

Anonymous:1990:BMa

[18] Anonymous. Back matter. Philosophy of Science, 57(1):177–182, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187633.

Anonymous:1990:FMa

[19] Anonymous. Front matter. Philosophy of Science, 57(1):??, March 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187616.

Thagard:1990:CSC

[20] Paul Thagard. The conceptual structure of the chemical revolution. Philosophy of Science, 57(2):183–209, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187831.

Shanahan:1990:EPS

[21] Timothy Shanahan. Evolution, phenotypic selection, and the units of selection. Philosophy of Science, 57(2):210–225, June 1990. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187832.

Cling:1990:DK

[22] Andrew D. Cling. Disappearance and knowledge. Philosophy of Science, 57(2):226–247, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187833.

Costa:1990:MTA

[23] Newton C. A. Da Costa and Steven French. The model-theoretic approach in the philosophy of science. Philosophy of Science, 57(2):248–265, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187834.

Osherson:1990:ASH

[24] Daniel N. Osherson and Scott Weinstein. On advancing simple hypotheses. Philosophy of Science, 57(2):266–277, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187835.

Mura:1990:WPS

[25] Alberto Mura. When probabilistic support is inductive. Philosophy of Science, 57(2):278–289, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187836.

Kary:1990:OCM

[26] Carla E. Kary. One causal mechanism in evolution: One unit of selection. Philosophy of Science, 57(2):290–296, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187837.

Ross:1990:APC

[27] Don Ross. Against positing central systems in the mind. Philosophy of Science, 57(2):297–312, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187838.

Brown:1990:DPR

[28] Harvey R. Brown. Does the Principle of Relativity imply Winnie’s (1970) Equal Passage Times Principle? Philosophy of Science, 57(2):313–324, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187839.

Demopoulos:1990:BRB

[29] William Demopoulos. Book review: Representation and Reality by Hilary Putnam. Philosophy of Science, 57(2):325–333, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187840.

Stuewer:1990:BRB

[30] Roger H. Stuewer. Book review: Selected Scientific Papers of Alfred Landé by A. O. Barut; A. van der Merwe. Philosophy of Science, 57(2):334– 335, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187841.

Solomon:1990:BRB

[31] Graham Solomon. Book review: Orthogonality and Spacetime Geometry by Robert Goldblatt. Philosophy of Science, 57(2):335–336, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187842.

Forster:1990:BRBa

[32] Malcolm R. Forster. Book review: Scientific Discovery: Computational Explorations of the Creative Process by Pat Langley; Herbert A. Simon; Gary L. Bradshaw; Jan M. Zytkow. Philosophy of Science, 57(2):336– 338, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187843.

Arntzenius:1990:BRB

[33] Frank Arntzenius. Book review: Probability and Causality by J. H. Fetzer. Philosophy of Science, 57(2):338–340, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187844.

Frankel:1990:BRB

[34] Henry Frankel. Book review: From Mineralogy to Geology: The Foundations of a Science, 1650–1830 by Rachel Laudan. Philosophy of Science, 57(2):340–342, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187845.

Clifton:1990:BRB

[35] Robert K. Clifton. Book review: Schrödinger: Centenary Celebration of a Polymath by C. W. Kilmister. Philosophy of Science, 57(2):342–343, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187846.

Gutting:1990:BRB

[36] Gary Gutting. Book review: Bachelard: Science and Objectivity by Mary Tiles. Philosophy of Science, 57(2):343–344, June 1990. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187847.

Bub:1990:BRB

[37] Jeffrey Bub. Book review: Niels Bohr’s Philosophy of Physics by Dugald Murdoch. Philosophy of Science, 57(2):344–347, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187848.

OBrien:1990:BRB

[38] David P. O’Brien. Book review: A Border Dispute: The Place of Logic in Psychology by John Macnamara. Philosophy of Science, 57(2):347– 349, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187849.

Kroes:1990:BRB

[39] Peter Kroes. Book review: An Architectonic for Science by Wolfgang Balzer; C. Ulises Moulines; Joseph D. Sneed. Philosophy of Science, 57 (2):349–350, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187850.

Anonymous:1990:BR

[40] Anonymous. Books received. Philosophy of Science, 57(2):351–356, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187851.

Anonymous:1990:BMb

[41] Anonymous. Back matter. Philosophy of Science, 57(2):357–368, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187852.

Anonymous:1990:FMb

[42] Anonymous. Front matter. Philosophy of Science, 57(2):??, June 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187830.

Carrier:1990:CCP

[43] Martin Carrier. Constructing or completing physical geometry? On the relation between theory and evidence in accounts of space-time structure. Philosophy of Science, 57(3):369–394, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187635.

Batterman:1990:ISM

[44] Robert W. Batterman. Irreversibility and statistical mechanics: A new approach? Philosophy of Science, 57(3):395–419, September 1990. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187636.

Kronz:1990:HLC

[45] Frederick M. Kronz. Hidden locality, conspiracy and superluminal signals. Philosophy of Science, 57(3):420–444, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187637.

Maloney:1990:MM

[46] J. Christopher Maloney. Mental misrepresentation. Philosophy of Science, 57(3):445–458, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187638.

Recker:1990:TMT

[47] Doren Recker. There’s more than one way to recognize a Darwinian: Lyell’s Darwinism. Philosophy of Science, 57(3):459–478, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187639.

Maher:1990:SAV

[48] Patrick Maher. Symptomatic acts and the value of evidence in causal decision theory. Philosophy of Science, 57(3):479–498, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187640.

Ettinger:1990:AOF

[49] Lia Ettinger, Eva Jablonka, and Peter McLaughlin. On the adaptations of organisms and the fitness of types. Philosophy of Science, 57(3):499–513, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187641.

Grimes:1990:TCH

[50] Thomas R. Grimes. Truth, content, and the hypothetico-deductive method. Philosophy of Science, 57(3):514–522, September 1990. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187642.

Heathcote:1990:UOI

[51] Adrian Heathcote. Unbounded operators and the incompleteness of quantum mechanics. Philosophy of Science, 57(3):523–534, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187643.

Good:1990:SFP

[52] I. J. Good. A suspicious feature of the Popper/Miller argument. Philosophy of Science, 57(3):535–536, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187644.

Davies:1990:BRB

[53] David Davies. Book review: Relativism Refuted by Harvey Siegel. Philosophy of Science, 57(3):537–539, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187645.

Recker:1990:BRB

[54] Doren Recker. Book review: Integrating Scientific Disciplines by William Bechtel. Philosophy of Science, 57(3):539–540, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187646.

Jacquette:1990:BRB

[55] Dale Jacquette. Book review: Parts: A Study in Ontology by Peter Simons. Philosophy of Science, 57(3):540–542, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187647.

Whitt:1990:BRB

[56] L. A. Whitt. Book review: Discovering Reality: Feminist Perspectives on Epistemology, Metaphysics, Methodology, and Philosophy of Science by Sandra Harding; Merrill B. Hintikka. Philosophy of Science, 57(3): 542–546, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187648.

Wylie:1990:BRB

[57] Alison Wylie. Book review: The Amateur and the Professional: Antiquarians, Historians and Archaeologists in Victorian England 1838–1886 by Philippa Levine; Science Encounters the Indian, 1820–1880: The Early Years of American Ethnology by Robert E. Bieder. Philosophy of Science, 57(3):546–548, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187649.

Anonymous:1990:BMc

[58] Anonymous. Back matter. Philosophy of Science, 57(3):549–552, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187650.

Anonymous:1990:FMc

[59] Anonymous. Front matter. Philosophy of Science, 57(3):??, September 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187634.

Greenwood:1990:TDN

[60] John D. Greenwood. Two dogmas of neo-empiricism: The “theoryinformity” of observation and the Quine–Duhem Thesis. Philosophy of Science, 57(4):553–574, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/187761. See reply [183].

Kincaid:1990:MBU

[61] Harold Kincaid. Molecular biology and the unity of science. Philosophy of Science, 57(4):575–593, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187762.

Teller:1990:PPI

[62] Paul Teller. Prolegomenon to a proper interpretation of quantum field theory. Philosophy of Science, 57(4):594–618, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187763.

Seigfried:1990:AQP

[63] Hans Seigfried. Autonomy and quantum physics: Nietzsche, Heidegger, and Heisenberg. Philosophy of Science, 57(4):619–630, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187764.

Dickman:1990:TQA

[64] Joel Dickman. Two qualms about functionalist Marxism. Philosophy of Science, 57(4):631–643, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187765.

Christensen:1990:IB

[65] David Christensen. The irrelevance of bootstrapping. Philosophy of Science, 57(4):644–662, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187766.

Smith:1990:QCB

[66] Gerrit Smith and Robert Weingard. Quantum cosmology and the beginning of the universe. Philosophy of Science, 57(4):663–667, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187767.

Richardson:1990:TAV

[67] Robert C. Richardson. The “tally argument” and the validation of psychoanalysis. Philosophy of Science, 57(4):668–676, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187768.

Sintonen:1990:DLS

[68] Matti Sintonen. Darwin’s long and short arguments. Philosophy of Science, 57(4):677–689, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187769.

Dupre:1990:PCR

[69] John Dupré. Probabilistic causality: A rejoinder to Ellery Eells. Philosophy of Science, 57(4):690–698, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187770. See [1].

Gemes:1990:HHH

[70] Ken Gemes. Horwich, Hempel, and hypothetico-deductivism. Philosophy of Science, 57(4):699–702, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187771.

Hinderliter:1990:MRH

[71] Hilton Hinderliter. More on Russell’s Hypothesis. Philosophy of Science, 57(4):703–711, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187772.

DiSalle:1990:BRB

[72] Robert DiSalle. Book review: Mach I, Mach II, Einstein, und Die Relativitätstheorie. Eine Fälschung und Ihre Folgen by Gereon Wolter. Philosophy of Science, 57(4):712–723, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187773.

Wians:1990:BRB

[73] William Wians. Book review: Aristotle on Nature and Living Things: Philosophical and Historical Studies. Presented to David M. Balme on His Seventieth Birthday by Allan Gotthelf. Philosophy of Science, 57(4): 724–725, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187774.

Ereshefsky:1990:BRB

[74] Marc Ereshefsky. Book review: Toward a New Philosophy of Biology by Ernst Mayr. Philosophy of Science, 57(4):725–727, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187775.

Forster:1990:BRBb

[75] Paul D. Forster. Book review: The Fortunes of Inquiry by Nicholas Jardine. Philosophy of Science, 57(4):727–729, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187776.

Teller:1990:BRB

[76] Paul Teller. Book review: Explaining Science: A Cognitive Approach by Ronald N. Giere. Philosophy of Science, 57(4):729–731, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187777.

Anonymous:1990:BMd

[77] Anonymous. Back matter. Philosophy of Science, 57(4):733–753, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187778.

Anonymous:1990:VI

[78] Anonymous. Volume information. Philosophy of Science, 57(4):735–740, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187759.

Anonymous:1990:FMd

[79] Anonymous. Front matter. Philosophy of Science, 57(4):??, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187760.

Horwich:1991:NNT

[80] Paul Horwich. On the nature and norms of theoretical commitment. Philosophy of Science, 58(1):1–14, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187886.

Patterson:1991:ISD

[81] Sarah Patterson. Individualism and semantic development. Philosophy of Science, 58(1):15–35, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187887.

Halpin:1991:WLF

[82] John F. Halpin. What is the logical form of probability assignment in quantum mechanics? Philosophy of Science, 58(1):36–60, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187888.

Mormann:1991:HPS

[83] Thomas Mormann. Husserl’s philosophy of science and the semantic approach. Philosophy of Science, 58(1):61–83, March 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187889.

Ereshefsky:1991:SHT

[84] Marc Ereshefsky. Species, higher taxa, and the units of evolution. Philosophy of Science, 58(1):84–101, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187890.

Mundy:1991:EUR

[85] Brent Mundy. Embedding and uniqueness in relationist theories. Philosophy of Science, 58(1):102–124, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187891.

Urbach:1991:RMC

[86] Peter Urbach. A reply to Mayo’s criticisms of Urbach’s “randomization and the design of experiments”. Philosophy of Science, 58(1):125–128, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187892.

Forman:1991:BRB

[87] Paul Forman. Book review: Intellectual Mastery of Nature; Theoretical Physics from Ohm to Einstein. Vol. 1, The Torch of Mathematics, 1800– 1870; Vol. 2, The Now Mighty Theoretical Physics, 1870–1925 by Christa Jungnickel; Russell McCormmach. Philosophy of Science, 58(1):129–132, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187893.

Hellman:1991:BRB

[88] Geoffrey Hellman. Book review: The Statue within: An Autobiography by François Jacob; F. Philip. Philosophy of Science, 58(1):132, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187894.

Falkenstein:1991:BRB

[89] Lorne Falkenstein. Book review: Particles and Ideas: Bishop Berkeley’s Corpuscularian Philosophy by Gabriel Moked. Philosophy of Science, 58 (1):133–135, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187895.

Staum:1991:BRB

[90] Martin S. Staum. Book review: Between Literature and Science: The Rise of Sociology by Wolf Lepenies; R. J. Hollingdale. Philosophy of Science, 58(1):135–136, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187896.

McMillan:1991:BRB

[91] John McMillan. Book review: The Popperian Legacy in Economics: Papers Presented at a Symposium in Amsterdam, December 1985 by Neil de March. Philosophy of Science, 58(1):136–138, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187897.

Bechtel:1991:BRB

[92] William Bechtel. Book review: Science as a Process by David L. Hull. Philosophy of Science, 58(1):138–139, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187898.

Anonymous:1991:BMb

[93] Anonymous. Back matter. Philosophy of Science, 58(1):140–146, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187899.

Anonymous:1991:FMb

[94] Anonymous. Front matter. Philosophy of Science, 58(1):??, March 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187885.

Stein:1991:RTO

[95] Howard Stein. On Relativity Theory and openness of the future. Philosophy of Science, 58(2):147–167, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187456. See [241].

Neander:1991:FSE

[96] Karen Neander. Functions as selected effects: The conceptual Analyst’s defense. Philosophy of Science, 58(2):168–184, June 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187457.

Jones:1991:RAW

[97] Roger Jones. Realism about what? Philosophy of Science, 58(2):185– 202, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187458.

Schwartz:1991:REM

[98] Justin Schwartz. Reduction, elimination, and the mental. Philosophy of Science, 58(2):203–220, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187459.

Allen:1991:CAN

[99] Colin Allen and Marc D. Hauser. Concept attribution in nonhuman animals: Theoretical and methodological problems in ascribing complex mental processes. Philosophy of Science, 58(2):221–240, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187460.

Batterman:1991:RPD

[100] Robert W. Batterman. Randomness and probability in dynamical theories: On the proposals of the Prigogine School. Philosophy of Science, 58(2):241–263, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187461.

Reisch:1991:DKK

[101] George A. Reisch. Did Kuhn kill logical empiricism? Philosophy of Science, 58(2):264–277, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187462.

Foss:1991:SPM

[102] Jeff Foss. On saving the phenomena and the mice: A reply to Bourgeois concerning van Fraassen’s image of science. Philosophy of Science, 58(2): 278–287, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187463.

Fetzer:1991:BRB

[103] James H. Fetzer. Book reviews: Scientific Explanation by Philip Kitcher; Wesley C. Salmon; Four Decades of Scientific Explanation by Wesley C. Salmon. Philosophy of Science, 58(2):288–306, June 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187464. See correction [121].

Schlagel:1991:CNF

[104] Richard H. Schlagel. Critical notice: Fine’s The Shaky Game (And Why NOA Is No Ark for Science). Philosophy of Science, 58(2):307–323, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187465.

Hooker:1991:BRB

[105] C. A. Hooker. Book review: Philosophical Foundations of Quantum Field Theory by Harvey R. Brown; Rom Harré. Philosophy of Science, 58(2): 324–329, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187466.

McGilvray:1991:BRB

[106] James A. McGilvray. Book review: Color for Philosophers by C. L. Hardin. Philosophy of Science, 58(2):329–331, June 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187467.

Anonymous:1991:BMc

[107] Anonymous. Back matter. Philosophy of Science, 58(2):332–336, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187468.

Anonymous:1991:FMc

[108] Anonymous. Front matter. Philosophy of Science, 58(2):??, June 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187455.

Cushing:1991:QTE

[109] James T. Cushing. Quantum theory and explanatory discourse: Endgame for understanding? Philosophy of Science, 58(3):337–358, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187936.

Martinez:1991:LRD

[110] Sergio Martinez. Lüders’s Rule as a description of individual state transformations. Philosophy of Science, 58(3):359–376, September 1991. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187937.

Deltete:1991:EE

[111] Robert Deltete and Reed Guy. Einstein and EPR. Philosophy of Science, 58(3):377–397, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187938.

Sober:1991:TAI

[112] Elliott Sober. Temporally asymmetric inference in a Markov process. Philosophy of Science, 58(3):398–410, September 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187939.

Smith:1991:NTR

[113] Quentin Smith. The new theory of reference entails absolute time and space. Philosophy of Science, 58(3):411–416, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187940.

Walton:1991:USB

[114] David Walton. The units of selection and the bases of selection. Philosophy of Science, 58(3):417–435, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187941.

Hahlweg:1991:NEP

[115] Kai Hahlweg. On the notion of evolutionary progress. Philosophy of Science, 58(3):436–451, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187942.

Miller:1991:CQC

[116] Steven I. Miller and Marcel Fredericks. A case for “qualitative confirmation” for the social and behavioral sciences. Philosophy of Science, 58(3): 452–467, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187943.

Enfield:1991:RES

[117] Patrick Enfield. Realism, empiricism and scientific revolutions. Philosophy of Science, 58(3):468–485, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187944.

Kukla:1991:CRP

[118] Andre Kukla. Criteria of rationality and the problem of logical sloth. Philosophy of Science, 58(3):486–490, September 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187945.

Niiniluoto:1991:GOT

[119] Ilkka Niiniluoto. Goldstick and O’Neill on “truer than”. Philosophy of Science, 58(3):491–495, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187946.

Gilman:1991:NO

[120] Daniel Gilman. The neurobiology of observation. Philosophy of Science, 58 (3):496–502, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187947.

Anonymous:1991:CSE

[121] Anonymous. Correction: Scientific explanation and four decades of scientific explan. Philosophy of Science, 58(3):502, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187948. See [103].

Healey:1991:BRB

[122] Richard A. Healey. Book review: Incompleteness, Nonlocality and Realism: A Prolegomenon to the Philosophy of Quantum Mechanics by Michael Redhead. Philosophy of Science, 58(3):503–505, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187949.

Morrison:1991:BRB

[123] Margaret Morrison. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Philosophy of Science, 58 (3):505–507, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187950.

Restivo:1991:BRB

[124] Sal Restivo. Book review: Beamtimes and Lifetimes: The World of High Energy Physics by Sharon Traweek. Philosophy of Science, 58(3):507–509, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187951.

Vicedo:1991:BRB

[125] Marga Vicedo. Book review: Philosophy of Biology Today by Michael Ruse. Philosophy of Science, 58(3):509, September 1991. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187952.

Anonymous:1991:BR

[126] Anonymous. Books received. Philosophy of Science, 58(3):510–519, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187953.

Anonymous:1991:BMd

[127] Anonymous. Back matter. Philosophy of Science, 58(3):520–522, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187954.

Anonymous:1991:FMd

[128] Anonymous. Front matter. Philosophy of Science, 58(3):??, September 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187935.

Mayo:1991:NES

[129] Deborah G. Mayo. Novel evidence and severe tests. Philosophy of Science, 58(4):523–552, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188479.

Waters:1991:TRA

[130] C. Kenneth Waters. Tempered realism about the force of selection. Philosophy of Science, 58(4):553–573, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188480.

Howson:1991:MMB

[131] Colin Howson and Allan Franklin. Maher, Mendeleev and Bayesianism. Philosophy of Science, 58(4):574–585, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188481.

Henderson:1991:TPG

[132] David K. Henderson. On the testability of psychological generalizations (psychological testability). Philosophy of Science, 58(4):586–606, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188482.

Rouse:1991:PPP

[133] Joseph Rouse. The politics of postmodern philosophy of science. Philosophy of Science, 58(4):607–627, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188483.

Redei:1991:BIR

[134] Miklos Redei. Bell’s inequalities, relativistic quantum field theory and the problem of hidden variables. Philosophy of Science, 58(4):628–638, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188484.

Verstraeten:1991:SCR

[135] Guido Verstraeten. Some critical remarks concerning Prigogine’s conception of temporal irreversibility. Philosophy of Science, 58(4):639–654, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188485.

Shapere:1991:LE

[136] Dudley Shapere. Leplin on essentialism. Philosophy of Science, 58(4): 655–677, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188486.

Nersessian:1991:MMR

[137] Nancy J. Nersessian. The method to “meaning”: A reply to Leplin. Philosophy of Science, 58(4):678–686, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188487.

Lipton:1991:CEC

[138] Peter Lipton. Contrastive explanation and causal triangulation. Philosophy of Science, 58(4):687–697, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188488.

Mundy:1991:BRB

[139] Brent Mundy. Book review: The Limitations of Deductivism by A. Grünbaum; W. Salmon. Philosophy of Science, 58(4):698–699, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188489.

Anonymous:1991:BMa

[140] Anonymous. Back matter. Philosophy of Science, 58(4):701–712, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188490.

Anonymous:1991:VI

[141] Anonymous. Volume information. Philosophy of Science, 58(4):703–707, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188477.

Anonymous:1991:FMa

[142] Anonymous. Front matter. Philosophy of Science, 58(4):??, December 1991. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188478.

Kronz:1992:PPT

[143] Frederick M. Kronz. The projection postulate and the time-energy uncertainty relation. Philosophy of Science, 59(1):1–15, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188115.

Elby:1992:SWE

[144] Andrew Elby. Should we explain the EPR correlations causally? Philosophy of Science, 59(1):16–25, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188116.

Butterfield:1992:DLM

[145] Jeremy Butterfield. David Lewis meets John Bell. Philosophy of Science, 59(1):26–43, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188117.

ODonnell:1992:KWA

[146] Rod O’Donnell. Keynes’s weight of argument and Popper’s paradox of ideal evidence. Philosophy of Science, 59(1):44–52, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188118.

Weston:1992:ATS

[147] Thomas Weston. Approximate truth and scientific realism. Philosophy of Science, 59(1):53–74, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188119.

Derksen:1992:DTA

[148] A. A. Derksen. Does the tally argument make Freud a sophisticated methodologist? Philosophy of Science, 59(1):75–101, March 1992. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188120.

Tuchanska:1992:WES

[149] Barbara Tuchanska. What is explained in science? Philosophy of Science, 59(1):102–119, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188121.

Maher:1992:DR

[150] Patrick Maher. Diachronic rationality. Philosophy of Science, 59(1):120– 141, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188122.

Sober:1992:SUS

[151] Elliott Sober. Screening-off and the units of selection. Philosophy of Science, 59(1):142–152, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188123.

Burian:1992:BRB

[152] Richard M. Burian. Book review: The Structure and Confirmation of Evolutionary Theory by Elisabeth A. Lloyd. Philosophy of Science, 59 (1):153–158, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188124.

Anonymous:1992:BMa

[153] Anonymous. Back matter. Philosophy of Science, 59(1):159–162, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188125.

Anonymous:1992:FMa

[154] Anonymous. Front matter. Philosophy of Science, 59(1):??, March 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188114.

vanFraassen:1992:GOJ

[155] Bas C. van Fraassen. The geometry of opinion: Jeffrey Shifts and linear operators. Philosophy of Science, 59(2):163–175, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188240.

Tieszen:1992:KGP

[156] Richard Tieszen. Kurt Gödel and phenomenology. Philosophy of Science, 59(2):176–194, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188241.

Dowe:1992:WSP

[157] Phil Dowe. Wesley Salmon’s process theory of causality and the conserved quantity theory. Philosophy of Science, 59(2):195–216, June 1992. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188242.

Bickle:1992:MAN

[158] John Bickle. Mental anomaly and the new mind-brain reductionism. Philosophy of Science, 59(2):217–230, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188243.

Trout:1992:TCM

[159] J. D. Trout. Theory-conjunction and mercenary reliance. Philosophy of Science, 59(2):231–245, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188244.

Fetzer:1992:WWS

[160] James H. Fetzer. What’s wrong with Salmon’s history: The third decade. Philosophy of Science, 59(2):246–262, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188245.

Winnie:1992:CC

[161] John A. Winnie. Computable chaos. Philosophy of Science, 59(2):263– 275, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188246.

Brandon:1992:STE

[162] Robert N. Brandon. A simple theory of evolution by natural selection. Philosophy of Science, 59(2):276–281, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188247.

Healey:1992:CRE

[163] Richard A. Healey. Causation, robustness, and EPR. Philosophy of Science, 59(2):282–292, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188248.

Linden:1992:SO

[164] Toby Linden. Shapere on observation. Philosophy of Science, 59(2):293– 299, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188249.

Malament:1992:BRB

[165] David B. Malament. Book review: Quantum Probability — Quantum Logic by Itamar Pitowsky. Philosophy of Science, 59(2):300–320, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188250.

Field:1992:BRB

[166] Hartry Field. Book review: Truth by Paul Horwich. Philosophy of Science, 59(2):321–330, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188251.

Cohen:1992:BRB

[167] Daniel H. Cohen. Book review: If P, Then Q: Conditionals and the Foundations of Reasoning by David H. Sanford. Philosophy of Science, 59(2): 331–332, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188252.

Ball:1992:BRB

[168] Stephen W. Ball. Book review: A Useful Inheritance: Evolutionary Aspects of the Theory of Knowledge by Nicholas Rescher. Philosophy of Science, 59(2):332–334, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188253.

Anonymous:1992:BMb

[169] Anonymous. Back matter. Philosophy of Science, 59(2):335–340, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188254.

Anonymous:1992:FMb

[170] Anonymous. Front matter. Philosophy of Science, 59(2):??, June 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188239.

Clendinnen:1992:NDC

[171] F. John Clendinnen. Nomic dependence and causation. Philosophy of Science, 59(3):341–360, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188152.

Narlikar:1992:CBC

[172] Jayant V. Narlikar. The concepts of “beginning” and “creation” in cosmology. Philosophy of Science, 59(3):361–371, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188153.

Leblanc:1992:PFM

[173] Hugues Leblanc and Peter Roeper. Probability functions: The matter of their recursive definability. Philosophy of Science, 59(3):372–388, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188154.

Neapolitan:1992:LFA

[174] Richard E. Neapolitan. A limiting frequency approach to probability based on the Weak Law of Large Numbers. Philosophy of Science, 59(3):389–407, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188155.

Hardcastle:1992:REE

[175] Valerie Gray Hardcastle. Reduction, explanatory extension, and the mind/brain sciences. Philosophy of Science, 59(3):408–428, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188156.

Tuomi:1992:ESS

[176] Juha Tuomi. Evolutionary synthesis: A search for the strategy. Philosophy of Science, 59(3):429–438, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188157.

Solomon:1992:SRH

[177] Miriam Solomon. Scientific rationality and human reasoning. Philosophy of Science, 59(3):439–455, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188158.

Stump:1992:NPS

[178] David Stump. Naturalized philosophy of science with a plurality of methods. Philosophy of Science, 59(3):456–460, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188159.

Hodge:1992:DAO

[179] M. J. S. Hodge. Darwin’s argument in the origin. Philosophy of Science, 59 (3):461–464, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188160.

Robinson:1992:AAR

[180] Joseph D. Robinson. Aims and achievements of the reductionist approach in biochemistry/molecular biology/cell biology: A response to Kincaid.

Philosophy of Science, 59(3):465–470, September 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188161.

Carroll:1992:UTP

[181] John W. Carroll. The unanimity theory and probabilistic sufficiency. Philosophy of Science, 59(3):471–479, September 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188162.

Wright:1992:GSH

[182] Edmond Wright. Gestalt switching: Hanson, Aronson, and Harre. Philosophy of Science, 59(3):480–486, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188163.

Klee:1992:DQD

[183] Robert Klee. In defense of the Quine–Duhem Thesis: A reply to Greenwood. Philosophy of Science, 59(3):487–491, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188164. See [60].

Kukla:1992:CI

[184] Andre Kukla. On the coherence of instrumentalism. Philosophy of Science, 59(3):492–497, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188165.

Resnik:1992:MRH

[185] David B. Resnik. Are methodological rules hypothetical imperatives? Philosophy of Science, 59(3):498–507, September 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188166.

Rynasiewicz:1992:WNT

[186] Robert Rynasiewicz. Why the new theory of reference does not entail absolute time and space. Philosophy of Science, 59(3):508–509, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188167.

Hacking:1992:BRBa

[187] Ian Hacking. Book review: Science in Action: How to Follow Scientists and Engineers through Society. The Pasteurization of France by Bruno Latour; Alan Sheridan; John Law. Philosophy of Science, 59(3):510–512, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188168.

Hunter:1992:BRB

[188] Daniel Hunter. Book reviews: Causation, Chance, and Credence: Proceedings of the Irvine Conference on Probability and Causation, Volume 1 by Brian Skyrms; William L. Harper; Causation in Decision, Belief Change, and Statistics: Proceedings of the Irvine Conference on Probability and Causation, Volume 2 by William L. Harper; Brian Skyrms. Philosophy of Science, 59(3):512–514, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188169.

Anonymous:1992:BR

[189] Anonymous. Books received. Philosophy of Science, 59(3):515–519, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188170.

Anonymous:1992:BMc

[190] Anonymous. Back matter. Philosophy of Science, 59(3):520–526, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188171.

Anonymous:1992:FMc

[191] Anonymous. Front matter. Philosophy of Science, 59(3):??, September 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188151.

Rosenkrantz:1992:JI

[192] R. D. Rosenkrantz. The justification of induction. Philosophy of Science, 59(4):527–539, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188128.

Christensen:1992:CHB

[193] David Christensen. Confirmational holism and Bayesian epistemology. Philosophy of Science, 59(4):540–557, December 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188129.

Barnes:1992:EUP

[194] Eric Barnes. Explanatory unification and the problem of asymmetry. Philosophy of Science, 59(4):558–571, December 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188130.

Rynasiewicz:1992:RHS

[195] Robert Rynasiewicz. Rings, holes and substantivalism: On the program of Leibniz algebras. Philosophy of Science, 59(4):572–589, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188131.

Bogen:1992:OTE

[196] Jim Bogen and Jim Woodward. Observations, theories and the evolution of the human spirit. Philosophy of Science, 59(4):590–611, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188132.

Whitt:1992:ITP

[197] Laurie Anne Whitt. Indices of theory promise. Philosophy of Science, 59 (4):612–634, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188133.

Enc:1992:FGD

[198] Berent Enc and Fred Adams. Functions and goal directedness. Philosophy of Science, 59(4):635–654, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188134.

Gasper:1992:RIG

[199] Philip Gasper. Reduction and instrumentalism in genetics. Philosophy of Science, 59(4):655–670, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188135.

Ereshefsky:1992:EP

[200] Marc Ereshefsky. Eliminative pluralism. Philosophy of Science, 59(4): 671–690, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188136.

Musgrave:1992:RAW

[201] Alan Musgrave. Realism about what? Philosophy of Science, 59(4): 691–697, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188137.

Goldstein:1992:BRB

[202] Bernard R. Goldstein. Book review: The General History of Astronomy. Vol. 2: Planetary Astronomy from the Renaissance to the Rise of Astrophysics. Part A: Tycho Brahe to Newton by Rene Taton; Curtis Wilson. Philosophy of Science, 59(4):698–700, December 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188138.

Jesseph:1992:BRB

[203] Douglas Jesseph. Book review: The Development of Newtonian Calculus in Britain 1700–1800 by Noccolo Guicciardini. Philosophy of Science, 59 (4):700–701, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188139.

Patrik:1992:BRB

[204] Linda E. Patrik. Book review: Critical Traditions in Contemporary Archaeology: Essays in the Philosophy, History and Socio-Politics of Archaeology by Valerie Pinsky; Alison Wylie. Philosophy of Science, 59(4): 701–703, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188140.

Sabloff:1992:BRB

[205] Jeremy A. Sabloff. Book review: A History of Archaeological Thought by Bruce G. Trigger. Philosophy of Science, 59(4):703–705, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188141.

Hacking:1992:BRBb

[206] Ian Hacking. Book reviews: The Uses of Experiment: Studies in the Natural Sciences by David Gooding; Trevor Pinch; Simon Schaffer; Experiment, Right or Wrong by Allan Franklin. Philosophy of Science, 59 (4):705–708, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188142.

Schabas:1992:BRB

[207] Margaret Schabas. Book review: More Heat than Light: Economics as Social Physics, Physics as Nature’s Economics by Philip Mirowski. Philosophy of Science, 59(4):708–710, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188143.

Gale:1992:BRB

[208] George Gale. Book review: An Intimate Relation: Studies in the History and Philosophy of Science by James R. Brown; Jurgen Mittelstrass. Philosophy of Science, 59(4):711–712, December 1992. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188144.

Shapin:1992:BRB

[209] Steven Shapin. Book review: The Rational and the Social by James Robert Brown. Philosophy of Science, 59(4):712–713, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188145.

Leplin:1992:BRB

[210] Jarrett Leplin. Book review: Science and Relativism: Some Key Controversies in the Philosophy of Science by Larry Laudan. Philosophy of Science, 59(4):713–714, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188146.

Lugg:1992:BRB

[211] Andrew Lugg. Book review: How Is Language Possible? Philosophical Reflections on the Evolution of Language and Knowledge by J. N. Hattiangadi. Philosophy of Science, 59(4):715–716, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188147.

Maudlin:1992:BRB

[212] Tim Maudlin. Book review: A Combinatorial Theory of Possibility by D. M. Armstrong. Philosophy of Science, 59(4):716–718, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188148.

Redhead:1992:BRB

[213] M. L. G. Redhead. Book review: World Enough and Space-Time: Absolute versus Relational Theories of Space and Time by John Earman. Philosophy of Science, 59(4):718–722, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188149.

Anonymous:1992:BMd

[214] Anonymous. Back matter. Philosophy of Science, 59(4):723–732, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188150.

Anonymous:1992:VI

[215] Anonymous. Volume information. Philosophy of Science, 59(4):725–730, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188126.

Anonymous:1992:FMd

[216] Anonymous. Front matter. Philosophy of Science, 59(4):??, December 1992. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188127.

Liston:1993:RMP

[217] Michael Liston. Reliability in mathematical physics. Philosophy of Science, 60(1):1–21, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188452.

Earman:1993:FDS

[218] John Earman and John D. Norton. Forever is a day: Supertasks in Pitowsky and Malament–Hogarth spacetimes. Philosophy of Science, 60 (1):22–42, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188453.

Batterman:1993:DC

[219] Robert W. Batterman. Defining chaos. Philosophy of Science, 60(1):43– 66, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188454.

Lahav:1993:WNT

[220] Ran Lahav. What neuropsychology tells us about consciousness. Philosophy of Science, 60(1):67–85, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188455.

Morton:1993:SCE

[221] Peter Morton. Supervenience and computational explanation in vision theory. Philosophy of Science, 60(1):86–99, March 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188456.

Khalidi:1993:CNJ

[222] Muhammad Ali Khalidi. Carving nature at the joints. Philosophy of Science, 60(1):100–113, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188457.

Sobel:1993:BIA

[223] Jordan Howard Sobel. Backward-induction arguments: A paradox regained. Philosophy of Science, 60(1):114–133, March 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188458.

Shapere:1993:AA

[224] Dudley Shapere. Astronomy and antirealism. Philosophy of Science, 60 (1):134–150, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188459.

Leblanc:1993:GCP

[225] Hugues Leblanc and Peter Roeper. Getting the constraints on Popper’s probability functions right. Philosophy of Science, 60(1):151–157, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188460.

Fuller:1993:BRBa

[226] Steve Fuller. Book review: Knowledge and Social Imagery (Second Edition) by David Bloor. Philosophy of Science, 60(1):158–170, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188461.

Ruse:1993:BRB

[227] Michael Ruse. Book reviews: The Non-Darwinian Revolution: Reinterpreting a Historical Myth by Peter Bowler; The Mendelian Revolution: The Emergence of Hereditarian Concepts in Modern Science and Society by Peter J. Bowler. Philosophy of Science, 60(1):171–172, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188462.

Cummins:1993:BRB

[228] Robert Cummins. Book review: A Theory of Content and Other Essays by Jerry Fodor. Philosophy of Science, 60(1):172–174, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188463.

Schouls:1993:BRB

[229] Peter A. Schouls. Book review: Diderot’s Dream by Wilda Anderson. Philosophy of Science, 60(1):174–176, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188464.

Anonymous:1993:BMa

[230] Anonymous. Back matter. Philosophy of Science, 60(1):177–182, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188465.

Anonymous:1993:FMa

[231] Anonymous. Front matter. Philosophy of Science, 60(1):??, March 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188451.

Maudlin:1993:BWW

[232] Tim Maudlin. Buckets of water and waves of space: Why spacetime is Probably a substance. Philosophy of Science, 60(2):183–203, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188350.

Mash:1993:BNI

[233] Roy Mash. Big numbers and induction in the case for extraterrestrial intelligence. Philosophy of Science, 60(2):204–222, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188351.

Melander:1993:HEE

[234] Peter Melander. How not to explain the errors of the immune system. Philosophy of Science, 60(2):223–241, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188352.

Owens:1993:CCP

[235] Joseph Owens. Content, causation, and psychophysical supervenience. Philosophy of Science, 60(2):242–261, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188353.

Seager:1993:FTC

[236] William Seager. Fodor’s theory of content: Problems and objections. Philosophy of Science, 60(2):262–277, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188354.

Schwartz:1993:FEM

[237] Justin Schwartz. Functional explanation and metaphysical individualism. Philosophy of Science, 60(2):278–301, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188355.

Juhl:1993:BRS

[238] Cory Juhl. Bayesianism and reliable scientific inquiry. Philosophy of Science, 60(2):302–319, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188356.

Skyrms:1993:MDC

[239] Brian Skyrms. A mistake in dynamic coherence arguments? Philosophy of Science, 60(2):320–328, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188357.

Maher:1993:HFP

[240] Patrick Maher. Howson and Franklin on prediction. Philosophy of Science, 60(2):329–340, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188358.

Maxwell:1993:RTO

[241] Nicholas Maxwell. On Relativity Theory and openness of the future. Philosophy of Science, 60(2):341–348, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188359. See [95].

Cheyne:1993:REF

[242] Colin Cheyne. Reduction, elimination, and firewalking. Philosophy of Science, 60(2):349–357, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188360.

Creath:1993:BRB

[243] Richard Creath. Book review: Reflexive Epistemology: The Philosophical Legacy of Otto Neurath by Danilo Zolo; D. McKie. Philosophy of Science, 60(2):359–360, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188361.

Fuller:1993:BRBb

[244] Steve Fuller. Book review: Science as Social Knowledge: Values and Objectivity in Scientific Inquiry by Helen E. Longino. Philosophy of Science, 60(2):360–362, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188362.

Green:1993:BRB

[245] Edward J. Green. Book review: The Invisible Hand by Bruna Ingrao; Giorgio Israel. Philosophy of Science, 60(2):362–363, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188363.

Anonymous:1993:BMb

[246] Anonymous. Back matter. Philosophy of Science, 60(2):364–372, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188364.

Anonymous:1993:FMb

[247] Anonymous. Front matter. Philosophy of Science, 60(2):??, June 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188349.

Anonymous:1993:DS

[248] Anonymous. The dimensions of selection. Philosophy of Science, 60(3): 373–395, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188081.

Sesardic:1993:HC

[249] Neven Sesardic. Heritability and causality. Philosophy of Science, 60(3): 396–418, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188082.

Nelkin:1993:WC

[250] Norton Nelkin. What is consciousness? Philosophy of Science, 60(3): 419–434, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188083.

Hausman:1993:LCE

[251] Daniel M. Hausman. Linking causal and explanatory asymmetry. Philosophy of Science, 60(3):435–451, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188084.

Downes:1993:SNP

[252] Stephen M. Downes. Socializing naturalized philosophy of science. Philosophy of Science, 60(3):452–468, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188085.

McKee:1993:TSM

[253] Bill McKee. A test of the scientific method. Philosophy of Science, 60(3): 469–476, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188086.

Gemes:1993:HDC

[254] Ken Gemes. Hypothetico-deductivism, content, and the natural axiomatization of theories. Philosophy of Science, 60(3):477–487, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188087.

Sharp:1993:RFT

[255] W. David Sharp and Niall Shanks. The rise and fall of time-symmetrized quantum mechanics. Philosophy of Science, 60(3):488–499, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188088.

Kaiser:1993:PAC

[256] Matthias Kaiser. Philosophers adrift? Comments on the alleged disunity of method. Philosophy of Science, 60(3):500–512, September 1993. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188089.

Melnick:1993:BRB

[257] Arthur Melnick. Book review: Kant’s Transcendental Psychology by P. Kitcher. Philosophy of Science, 60(3):513–515, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188090.

Cartwright:1993:BRB

[258] Nancy Cartwright. Book review: The History of Econometric Ideas: Historical Perspectives on Modern Economics by M. S. Morgan. Philosophy of Science, 60(3):515–516, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188091.

Quinn:1993:BRB

[259] Philip L. Quinn. Book review: Theology for a Scientific Age: Being and Becoming — Natural and Divine by A. Peacocke. Philosophy of Science, 60 (3):516–518, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188092.

McGee:1993:BRB

[260] Vann McGee. Book review: A Theory of Counterfactuals by Igal Kvart. Philosophy of Science, 60(3):518–519, September 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188093.

Gale:1993:BRB

[261] George Gale. Book review: Universes by John Leslie. Philosophy of Science, 60(3):519–521, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188094.

Baltas:1993:BRBa

[262] Aristides Baltas. Book review: Moral Consciousness and Communicative Action by Jurgen Habermas; Christian Lenhardt; Shierry Weber Nicholsen. Philosophy of Science, 60(3):521–523, September 1993. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188095.

Anonymous:1993:BR

[263] Anonymous. Books received. Philosophy of Science, 60(3):524–528, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188096.

Anonymous:1993:BMc

[264] Anonymous. Back matter. Philosophy of Science, 60(3):529–533, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188097.

Anonymous:1993:FMc

[265] Anonymous. Front matter. Philosophy of Science, 60(3):??, September 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188080.

Gatens-Robinson:1993:WFW

[266] Eugenie Gatens-Robinson. Why falsification is the wrong paradigm for evolutionary epistemology: An analysis of Hull’s selection theory. Philosophy of Science, 60(4):535–557, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188186.

Sibelius:1993:MFW

[267] Patrick Sibelius. A major failure within modern analytic philosophy. Philosophy of Science, 60(4):558–567, December 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188187.

Peressini:1993:GEA

[268] Anthony Peressini. Generalizing evolutionary altruism. Philosophy of Science, 60(4):568–586, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188188.

Uebel:1993:NPS

[269] Thomas E. Uebel. Neurath’s protocol statements: A naturalistic theory of data and pragmatic theory of theory acceptance. Philosophy of Science, 60 (4):587–607, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188189.

Redei:1993:PSC

[270] Miklos Redei. Are prohibitions of superluminal causation by stochastic Einstein locality and by absence of Lewisian probabilistic counterfactual causality equivalent? Philosophy of Science, 60(4):608–618, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188190.

Liu:1993:ATQ

[271] Chuang Liu. The Arrow of time in quantum gravity. Philosophy of Science, 60(4):619–637, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188191.

Grunbaum:1993:NCB

[272] Adolf Grünbaum. Narlikar’s “creation” of the big Bang universe was a mere origination. Philosophy of Science, 60(4):638–646, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188192.

Baltas:1993:BRBb

[273] Aristides Baltas. Book review: Philosophy and the Spontaneous Philosophy of the Scientists and Other Essays by Louis Althusser. Philosophy of Science, 60(4):647–658, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188193.

Dieterle:1993:BRB

[274] Jill Dieterle and Stewart Shapiro. Book review: Realism in Mathematics by Penelope Maddy. Philosophy of Science, 60(4):659–660, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188194.

Smith:1993:BRB

[275] Joel M. Smith. Book review: Laws and Symmetry by Bas C. van Fraassen. Philosophy of Science, 60(4):661–662, December 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188195.

Hatfield:1993:BRB

[276] Gary Hatfield. Book review: Historical Roots of Cognitive Science: The Rise of a Cognitive Theory of Perception from Antiquity to the Nineteenth Century by Theo C. Meyering. Philosophy of Science, 60(4):662–666, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188196.

Cushing:1993:BRB

[277] James T. Cushing. Book review: Too Hot to Handle: The Race for Cold Fusion by Frank Close. Philosophy of Science, 60(4):666–667, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188197.

Leslie:1993:BRB

[278] John Leslie. Book review: Cosmos and Anthropos: A Philosophical Interpretation of the Anthropic Cosmological Principle by Errol E. Harris. Philosophy of Science, 60(4):667–669, December 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188198.

Perlman:1993:BRB

[279] Mark Perlman. Book review: Stabilizing Dynamics: Constructing Economic Knowledge by E. Roy Weintraub. Philosophy of Science, 60(4): 669–671, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188199.

Woodward:1993:BRB

[280] Jim Woodward. Book review: The Chances of Explanation: Causal Explanation in the Social, Medical and Physical Sciences by Paul Humphreys. Philosophy of Science, 60(4):671–673, December 1993. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188200.

Buchwald:1993:BRB

[281] Jed Z. Buchwald. Book review: Methodological Aspects of the Development of Low Temperature Physics 1881–1956: Concepts out of Context by K. Gavroglu; Y. Goudaroulis. Philosophy of Science, 60(4):673–675, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188201.

Peacocke:1993:BRB

[282] Christopher Peacocke. Book review: The Imagery Debate by Michael Tye. Philosophy of Science, 60(4):675–677, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188202.

Maffie:1993:BRB

[283] James Maffie. Book review: Scientism: Philosophy and the Infatuation with Science by Tom Sorell. Philosophy of Science, 60(4):677–679, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188203.

Hobbs:1993:BRB

[284] Jesse Hobbs. Book review: Inference to the Best Explanation by Peter Lipton. Philosophy of Science, 60(4):679–680, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188204.

Anonymous:1993:BMd

[285] Anonymous. Back matter. Philosophy of Science, 60(4):681–695, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188205.

Anonymous:1993:VI

[286] Anonymous. Volume information. Philosophy of Science, 60(4):683–688, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188184.

Anonymous:1993:FMd

[287] Anonymous. Front matter. Philosophy of Science, 60(4):??, December 1993. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188185.

Marinoff:1994:RBP

[288] Louis Marinoff. A resolution of Bertrand’s paradox. Philosophy of Science, 61(1):1–24, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188286.

Rothbart:1994:ES

[289] Daniel Rothbart and Suzanne W. Slayden. The epistemology of a spectrometer. Philosophy of Science, 61(1):25–38, March 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188287.

Rollins:1994:DPE

[290] Mark Rollins. Deep plasticity: The encoding approach to perceptual change. Philosophy of Science, 61(1):39–54, March 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188288.

Wilson:1994:CDT

[291] Robert A. Wilson. Causal depth, theoretical appropriateness, and individualism in psychology. Philosophy of Science, 61(1):55–75, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188289.

Horan:1994:SCE

[292] Barbara L. Horan. The statistical character of evolutionary theory. Philosophy of Science, 61(1):76–95, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188290.

Kvart:1994:CI

[293] Igal Kvart. Causal Independence. Philosophy of Science, 61(1):96–114, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188291.

Trenholme:1994:AS

[294] Russell Trenholme. Analog simulation. Philosophy of Science, 61(1):115– 131, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188292.

Churchman:1994:WPS

[295] C. West Churchman. What is philosophy of science? Philosophy of Science, 61(1):132–141, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188293.

Richardson:1994:BRB

[296] Alan Richardson. Book review: The Semantic Tradition from Kant to Carnap: To the Vienna Station by Alberto Coffa. Philosophy of Science, 61(1):142–144, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188294.

Turnbull:1994:BRB

[297] Robert G. Turnbull. Book review: The Chain of Change: A Study of Aristotle’s Physics VII by Robert Wardy. Philosophy of Science, 61(1): 144–145, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188295.

Pennock:1994:BRB

[298] Robert T. Pennock. Book review: Explaining Explanation by David-Hillel Ruben. Philosophy of Science, 61(1):146–147, March 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188296.

Townsend:1994:BRB

[299] Burke Townsend. Book review: Creative Understanding: Philosophical Reflections on Physics by Roberto Torretti. Philosophy of Science, 61 (1):147–148, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188297.

Cushing:1994:BRB

[300] James T. Cushing. Book review: Niels Bohr: His Heritage and Legacy by Jan Faye. Philosophy of Science, 61(1):149–150, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188298.

Kitcher:1994:BRB

[301] Patricia Kitcher. Book review: Darwin’s Influence on Freud: A Tale of Two Sciences by Lucille B. Ritvo. Philosophy of Science, 61(1):150– 151, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188299.

Pinkus:1994:BRB

[302] Rosa Lynn Pinkus. Book review: Embryo Experimentation by Peter Singer; Helga Kuhse; Stephen Buckle; Karen Dawson; Pascal Kasimba. Philosophy of Science, 61(1):151–153, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/188300.

Wilson:1994:BRB

[303] Bradley E. Wilson. Book review: Theory Change in Science: Strategies from Mendelian Genetics by Lindley Darden. Philosophy of Science, 61 (1):153–155, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188301.

Kronz:1994:BRB

[304] Frederick M. Kronz. Book review: Nature’s Capacities and Their Measurement by Nancy D. Cartwright. Philosophy of Science, 61(1):155–157, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188302.

Anonymous:1994:BMa

[305] Anonymous. Back matter. Philosophy of Science, 61(1):158–162, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188303.

Anonymous:1994:FMa

[306] Anonymous. Front matter. Philosophy of Science, 61(1):??, March 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188285.

Waters:1994:GMM

[307] C. Kenneth Waters. Genes made molecular. Philosophy of Science, 61(2): 163–185, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188207.

Ereshefsky:1994:SPL

[308] Marc Ereshefsky. Some problems with the Linnaean hierarchy. Philosophy of Science, 61(2):186–205, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188208.

Griffiths:1994:CCF

[309] P. E. Griffiths. Cladistic classification and functional explanation. Philosophy of Science, 61(2):206–227, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188209.

Shrader-Frechette:1994:AEL

[310] Kristin Shrader-Frechette and Earl D. Mccoy. Applied ecology and the logic of case studies. Philosophy of Science, 61(2):228–249, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188210.

Stone:1994:DBT

[311] Abraham D. Stone. Does the Bohm theory solve the measurement problem? Philosophy of Science, 61(2):250–266, June 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188211.

Zangari:1994:NTC

[312] Mark Zangari. A new twist in the conventionality of simultaneity debate. Philosophy of Science, 61(2):267–275, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188212.

Giere:1994:CSS

[313] Ronald N. Giere. The cognitive structure of scientific theories. Philosophy of Science, 61(2):276–296, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188213.

Salmon:1994:CC

[314] Wesley C. Salmon. Causality without counterfactuals. Philosophy of Science, 61(2):297–312, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188214.

Murad:1994:BRB

[315] Hazim Murad. Book reviews: German Science by Pierre Duhem; John Lyon; Pierre Duhem: Philosophy and History in the Work of a Believing Physicist by R. N. Martin. Philosophy of Science, 61(2):313–315, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188215.

Kincaid:1994:BRB

[316] Harold Kincaid. Book reviews: Economics — Mathematical Politics or Science of Diminishing Returns? by Alexander Rosenberg; Essays on Philosophy and Economic Methodology by Daniel Hausman. Philosophy of Science, 61(2):315–318, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188216.

Barrotta:1994:BRB

[317] Pierluigi Barrotta. Book review: The Inexact and Separate Science of Economics by Daniel M. Hausman. Philosophy of Science, 61(2):318– 319, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188217.

Averill:1994:BRB

[318] Edward Wilson Averill. Book review: Sensations: A Defense of Type Materialism by Christopher S. Hill. Philosophy of Science, 61(2):319– 321, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188218.

Buroker:1994:BRB

[319] Jill Vance Buroker. Book review: Kant and the Exact Sciences by Michael Friedman. Philosophy of Science, 61(2):321–322, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188219.

Anonymous:1994:BMb

[320] Anonymous. Back matter. Philosophy of Science, 61(2):323–328, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188220.

Anonymous:1994:FMb

[321] Anonymous. Front matter. Philosophy of Science, 61(2):??, June 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188206.

Achinstein:1994:SE

[322] Peter Achinstein. Stronger evidence. Philosophy of Science, 61(3):329– 350, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188049.

Brown:1994:RJB

[323] Harold I. Brown. Reason, judgement and Bayes’s law. Philosophy of Science, 61(3):351–369, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188050.

Huggett:1994:IQF

[324] Nick Huggett and Robert Weingard. Interpretations of quantum field theory. Philosophy of Science, 61(3):370–388, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188051.

Mormann:1994:AKL

[325] Thomas Mormann. Accessibility, kinds, and laws: A structural explication. Philosophy of Science, 61(3):389–406, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188052.

Wallis:1994:RII

[326] Charles Wallis. Representation and the imperfect ideal. Philosophy of Science, 61(3):407–428, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188053.

Quillian:1994:CIE

[327] M. Ross Quillian. A content-independent explanation of sciences effectiveness. Philosophy of Science, 61(3):429–448, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188054.

Bunzl:1994:SAR

[328] Martin Bunzl. Scientific abstraction and the Realist impulse. Philosophy of Science, 61(3):449–456, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188055.

Muller:1994:ALC

[329] F. A. Muller and Jeremy Butterfield. Is algebraic Lorentz-covariant quantum field theory stochastic Einstein local? Philosophy of Science, 61(3): 457–474, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188056.

Brandon:1994:SBS

[330] Robert N. Brandon, Janis Antonovics, Richard Burian, Scott Carson, Greg Cooper, Paul Sheldon Davies, Christopher Horvath, Brent D. Mishler, Robert C. Richardson, Kelly Smith, and Peter Thrall. Sober on Brandon on screening-off and the levels of selection. Philosophy of Science, 61 (3):475–486, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188057.

Maio:1994:BRB

[331] Maria Concetta Di Maio. Book review: Notes on Philosophy, Probability and Mathematics by Frank Plumpton Ramsey; Maria Carla Galavotti. Philosophy of Science, 61(3):487–489, September 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188058.

Mullen:1994:BRB

[332] John D. Mullen. Book review: Utility Theories: Measurements and Applications by Ward Edwards. Philosophy of Science, 61(3):489–490, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188059.

Anonymous:1994:BR

[333] Anonymous. Books received. Philosophy of Science, 61(3):491–495, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188060.

Anonymous:1994:BMc

[334] Anonymous. Back matter. Philosophy of Science, 61(3):496–501, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188061.

Anonymous:1994:FMc

[335] Anonymous. Front matter. Philosophy of Science, 61(3):??, September 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188048.

Skyrms:1994:DML

[336] Brian Skyrms. Darwin meets the logic of decision: Correlation in evolutionary game theory. Philosophy of Science, 61(4):503–528, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188332.

Mayr:1994:RFT

[337] Ernst Mayr. Reasons for the failure of theories. Philosophy of Science, 61 (4):529–533, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188333.

Sober:1994:CRP

[338] Elliott Sober and David Sloan Wilson. A critical review of philosophical work on the units of selection problem. Philosophy of Science, 61(4):

534–555, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188334.

Amundson:1994:TCC

[339] Ron Amundson. Two concepts of constraint: Adaptationism and the challenge from developmental biology. Philosophy of Science, 61(4):556– 578, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188335.

Wachbroit:1994:NBC

[340] Robert Wachbroit. Normality as a biological concept. Philosophy of Science, 61(4):579–591, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188336.

Hoefer:1994:EEU

[341] Carl Hoefer and Alexander Rosenberg. Empirical equivalence, underdetermination, and systems of the world. Philosophy of Science, 61(4):592–607, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188337.

Balashov:1994:DQM

[342] Yuri Balashov. Duhem, Quine, and the multiplicity of scientific tests. Philosophy of Science, 61(4):608–628, December 1994. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188338.

Thagard:1994:MSG

[343] Paul Thagard. Mind, society, and the growth of knowledge. Philosophy of Science, 61(4):629–645, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188339.

Pinnick:1994:FEI

[344] Cassandra L. Pinnick. Feminist epistemology: Implications for philosophy of science. Philosophy of Science, 61(4):646–657, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188340.

Saidel:1994:CCP

[345] Eric Saidel. Content and causal powers. Philosophy of Science, 61(4): 658–665, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188341.

Leplin:1994:BRB

[346] Jarrett Leplin. Book review: The Advancement of Science: Science without Legend, Objectivity without Illusion by Philip Kitcher. Philosophy of Science, 61(4):666–671, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188342.

Worrall:1994:BRB

[347] John Worrall. Book review: Bayes or Bust? A Critical Examination of Bayesian Confirmation Theory by John Earman. Philosophy of Science, 61 (4):672–673, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188343.

Lennox:1994:BRB

[348] James G. Lennox. Book review: The Meaning of Evolution: The Morphological Construction and Ideological Reconstruction of Darwin’s Theory by Robert J. Richards. Philosophy of Science, 61(4):673–675, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188344.

Lynch:1994:BRB

[349] Michael Lynch. Book review: Making Science: Between Nature and Society by Stephen Cole. Philosophy of Science, 61(4):675–677, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188345.

Hardcastle:1994:BRB

[350] Gary Hardcastle. Book review: The Olympian Dreams and Youthful Rebellion of Rene Descartes by John R. Cole. Philosophy of Science, 61(4): 677–679, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188346.

Risjord:1994:BRB

[351] Mark Risjord. Book review: Bloodsucking Witchcraft: An Epistemological Study of Anthropomorphic Supernaturalism in Rural Tlaxcala by Hugo G. Nutini; John M. Roberts. Philosophy of Science, 61(4):679–681, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188347.

Anonymous:1994:BMd

[352] Anonymous. Back matter. Philosophy of Science, 61(4):682–692, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188348.

Anonymous:1994:VI

[353] Anonymous. Volume information. Philosophy of Science, 61(4):683–688, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188330.

Anonymous:1994:FMd

[354] Anonymous. Front matter. Philosophy of Science, 61(4):??, December 1994. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188331.

Ramsey:1995:CR

[355] Jeffry L. Ramsey. Construction by reduction. Philosophy of Science, 62 (1):1–20, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188031.

Jones:1995:RUT

[356] Todd Jones. Reductionism and the unification theory of explanation. Philosophy of Science, 62(1):21–30, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188032.

Flichman:1995:HSA

[357] Eduardo H. Flichman. Hard and soft accidental uniformities. Philosophy of Science, 62(1):31–43, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188033.

Castelao-Lawless:1995:PHP

[358] Teresa Castelao-Lawless. Phenomenotechnique in historical perspective: Its origins and implications for philosophy of science. Philosophy of Science, 62(1):44–59, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188034.

Reiner:1995:HER

[359] Richard Reiner and Robert Pierson. Hacking’s experimental realism: An untenable middle ground. Philosophy of Science, 62(1):60–69, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188035.

Stanford:1995:PAR

[360] P. Kyle Stanford. For pluralism and against realism about species. Philosophy of Science, 62(1):70–91, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188036.

Kaye:1995:LT

[361] Lawrence J. Kaye. The languages of thought. Philosophy of Science, 62(1):92–110, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188037.

Wayne:1995:BDE

[362] Andrew Wayne. Bayesianism and diverse evidence. Philosophy of Science, 62(1):111–121, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188038.

See reply [520].

Dickson:1995:ERE

[363] Michael Dickson. An empirical reply to empiricism: Protective measurement opens the door for quantum realism. Philosophy of Science, 62(1): 122–140, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188039.

Reeder:1995:PPD

[364] Nick Reeder. Are physical properties dispositions? Philosophy of Science, 62(1):141–149, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188040.

Raftopoulos:1995:WCS

[365] Athanasse Raftopoulos. Was Cartesian science ever meant to be a priori? A comment on Hatfield. Philosophy of Science, 62(1):150–160, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188041.

Goldstein:1995:BRB

[366] Bernard R. Goldstein. Book review: New Astronomy by Johannes Kepler; William H. Donahue. Philosophy of Science, 62(1):161–162, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188042.

Manders:1995:BRB

[367] Kenneth L. Manders. Book review: The Magic of Numbers and Motion: The Scientific Career of Rene Descartes by William S. Shea. Philosophy of Science, 62(1):162–164, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188043.

Salmon:1995:BRB

[368] Wesley C. Salmon. Book review: The Ambiguous Frog: The Galvani–Volta Controversy on Animal Electricity by Marcello Pera; Jonathan Mandelbaum. Philosophy of Science, 62(1):164–166, March 1995. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188044.

Kitcher:1995:BRB

[369] Patricia Kitcher. Book review: Validation in the Clinical Theory of Psychoanalysis by Adolf Grünbaum. Philosophy of Science, 62(1):166–167, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188045.

Sulloway:1995:BRB

[370] Frank J. Sulloway. Book review: Freud’s Dream: A Complete Interdisciplinary Science of Mind by Patricia Kitcher. Philosophy of Science, 62 (1):168–170, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188046.

Anonymous:1995:BMa

[371] Anonymous. Back matter. Philosophy of Science, 62(1):171–178, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188047.

Anonymous:1995:FMa

[372] Anonymous. Front matter. Philosophy of Science, 62(1):??, March 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188030.

Egan:1995:FPC

[373] Frances Egan. Folk psychology and cognitive architecture. Philosophy of Science, 62(2):179–196, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188429.

Breuer:1995:IAS

[374] Thomas Breuer. The impossibility of accurate state self-measurements. Philosophy of Science, 62(2):197–214, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188430.

Barnes:1995:TTA

[375] Eric Barnes. Truthlikeness, translation, and approximate causal explanation. Philosophy of Science, 62(2):215–226, June 1995. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188431.

Schwartz:1995:MCI

[376] Robert Schwartz. Is mathematical competence innate? Philosophy of Science, 62(2):227–240, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188432.

Mitchell:1995:TDB

[377] Sam Mitchell. Toward a defensible bootstrapping. Philosophy of Science, 62(2):241–260, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188433.

Montgomery:1995:EEC

[378] Richard Montgomery. Explanation and evaluation in cognitive science. Philosophy of Science, 62(2):261–282, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188434.

Mulaik:1995:MOO

[379] Stanley A. Mulaik. The metaphoric origins of objectivity, subjectivity, and consciousness in the direct perception of reality. Philosophy of Science, 62(2):283–303, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188435.

Hitchcock:1995:SER

[380] Christopher Read Hitchcock. Salmon on explanatory relevance. Philosophy of Science, 62(2):304–320, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188436.

Dowe:1995:CCQ

[381] Phil Dowe. Causality and conserved quantities: A reply to Salmon. Philosophy of Science, 62(2):321–333, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188437.

Kitchener:1995:BRB

[382] Richard F. Kitchener. Book review: The Natural and the Normative by Gary Hatfield. Philosophy of Science, 62(2):334–335, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188438.

Ryckman:1995:BRB

[383] T. A. Ryckman. Book review: Overcoming Logical Positivism from within: The Emergence of Neurath’s Naturalism in the Vienna Circle’s Protocol Sentence Debate by Thomas E. Uebel. Philosophy of Science, 62(2):335– 337, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188439.

Rynasiewicz:1995:BRB

[384] Robert Rynasiewicz. Book review: Physics and Chance: Philosophical Issues in the Foundations of Statistical Mechanics by Lawrence Sklar. Philosophy of Science, 62(2):337–338, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188440.

Post:1995:BRB

[385] John F. Post. Book review: Supervenience and Mind: Selected Philosophical Essays by Jaegwon Kim. Philosophy of Science, 62(2):338–340, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188441.

Smith:1995:BRB

[386] Quentin Smith. Book review: Modalities: Philosophical Essays by Ruth Barcan Marcus. Philosophy of Science, 62(2):340–341, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188442.

Massey:1995:BRB

[387] Gerald J. Massey. Book review: The Laboratory of the Mind: Thought Experiments in the Natural Sciences by James Robert Brown. Philosophy of Science, 62(2):341–343, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188443.

Kyburg:1995:BRB

[388] Henry E. Kyburg, Jr. Book review: Betting on Theories by Patrick Maher. Philosophy of Science, 62(2):343–345, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188444.

Thagard:1995:BRB

[389] Paul Thagard. Book review: What Is Cognitive Science? by Barbara Von Eckardt. Philosophy of Science, 62(2):345–346, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188445.

Woody:1995:BRB

[390] Andrea Woody. Book review: Simians, Cyborgs, and Women: The Reinvention of Nature by Donna Haraway. Philosophy of Science, 62(2):346– 348, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188446.

MacKinnon:1995:BRB

[391] Edward MacKinnon. Book review: From c-Numbers to q-Numbers: The Classical History of Quantum Theory by Olivier Darrigol. Philosophy of Science, 62(2):348–350, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188447.

Tuchanska:1995:BRB

[392] Barbara Tuchanska. Book review: We Have Never Been Modern by Bruno Latour. Philosophy of Science, 62(2):350–351, June 1995. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188448.

Thalos:1995:BRB

[393] Mariam Thalos. Book review: The Disorder of Things: Metaphysical Foundations of the Disunity of Science by John Dupre. Philosophy of Science, 62(2):351–353, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188449.

Anonymous:1995:BMb

[394] Anonymous. Back matter. Philosophy of Science, 62(2):354–356, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188450.

Anonymous:1995:FMb

[395] Anonymous. Front matter. Philosophy of Science, 62(2):??, June 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188428.

Arntzenius:1995:HCC

[396] Frank Arntzenius. A heuristic for conceptual change. Philosophy of Science, 62(3):357–369, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188492.

Melnyk:1995:TCR

[397] Andrew Melnyk. Two cheers for reductionism: Or, the dim prospects for non-reductive materialism. Philosophy of Science, 62(3):370–388, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188493.

Falk:1995:WU

[398] Arthur Falk. Wisdom updated. Philosophy of Science, 62(3):389–403, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188494.

Hardcastle:1995:SOO

[399] Gary L. Hardcastle. S. S. Stevens and the origins of operationism. Philosophy of Science, 62(3):404–424, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188495.

Leeds:1995:HDA

[400] Stephen Leeds. Holes and determinism: Another look. Philosophy of Science, 62(3):425–437, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188496.

Culp:1995:OEI

[401] Sylvia Culp. Objectivity in experimental inquiry: Breaking datatechnique circles. Philosophy of Science, 62(3):438–458, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188497.

Seager:1995:GTV

[402] William Seager. Ground truth and virtual reality: Hacking vs. van Fraassen. Philosophy of Science, 62(3):459–478, September 1995. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188498.

Maudlin:1995:WBT

[403] Tim Maudlin. Why Bohm’s theory solves the measurement problem. Philosophy of Science, 62(3):479–483, September 1995. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188499.

Pauri:1995:BRB

[404] Massimo Pauri. Book review: Philosophical Problems of the Internal and External Worlds: Essays on the Philosophy of Adolf Grunbaum by John Earman; Allen I. Janis; Gerald J. Massey; Nicholas Rescher. Philosophy of Science, 62(3):484–487, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188500.

Hull:1995:BRB

[405] David L. Hull. Book review: The Golem: What Everyone Should Know about Science by Harry Collins; Trevor Pinch. Philosophy of Science, 62 (3):487–488, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188501.

Anonymous:1995:BMc

[406] Anonymous. Back matter. Philosophy of Science, 62(3):489–493, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188502.

Anonymous:1995:FMc

[407] Anonymous. Front matter. Philosophy of Science, 62(3):??, September 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188491.

Malament:1995:NCR

[408] David B. Malament. Is Newtonian cosmology really inconsistent? Philosophy of Science, 62(4):489–510, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188548.

Norton:1995:FNC

[409] John D. Norton. The force of Newtonian cosmology: Acceleration is relative. Philosophy of Science, 62(4):511–522, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188549.

Enc:1995:UB

[410] Berent Enc. Units of behavior. Philosophy of Science, 62(4):523–542, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188550.

Kennedy:1995:EFQ

[411] J. B. Kennedy. On the empirical foundations of the quantum no-signalling proofs. Philosophy of Science, 62(4):543–560, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188551.

Culler:1995:BBN

[412] Madison Culler. Beyond bootstrapping: A new account of evidential relevance. Philosophy of Science, 62(4):561–579, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188552.

Kosso:1995:CPT

[413] Peter Kosso and Cynthia Kosso. Central place theory and the reciprocity between theory and evidence. Philosophy of Science, 62(4):581–598, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188553.

Gunn:1995:RQM

[414] David Gunn and Indrakumar Vetharaniam. Relativistic quantum mechanics and the conventionality of simultaneity. Philosophy of Science, 62(4): 599–608, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188554.

Allen:1995:BFA

[415] Colin Allen and Marc Bekoff. Biological function, adaptation, and natural design. Philosophy of Science, 62(4):609–622, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188555.

Anonymous:1995:BMd

[416] Anonymous. Back matter. Philosophy of Science, 62(4):623–635, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188556.

Anonymous:1995:VI

[417] Anonymous. Volume information. Philosophy of Science, 62(4):625–630, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188546.

Anonymous:1995:FMd

[418] Anonymous. Front matter. Philosophy of Science, 62(4):??, December 1995. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188547.

Allchin:1996:PEW

[419] Douglas Allchin. Points East and west: Acupuncture and comparative philosophy of science. Philosophy of Science, 63:S107–S115, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188517.

Anonymous:1996:BMd

[420] Anonymous. Back matter. Philosophy of Science, 63:S347, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188545.

Anonymous:1996:FMd

[421] Anonymous. Front matter. Philosophy of Science, 63:i–iv, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188503.

Ariew:1996:IC

[422] Andre Ariew. Innateness and canalization. Philosophy of Science, 63: S19–S27, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188507.

Bandyopadhayay:1996:CFP

[423] Prasanta S. Bandyopadhayay, Robert J. Boik, and Prasun Basu. The curve fitting problem: A Bayesian approach. Philosophy of Science, 63: S264–S272, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188535.

Barrett:1996:OAB

[424] Jeffrey A. Barrett. Oracles, aesthetics, and Bayesian consensus. Philosophy of Science, 63:S273–S280, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188536.

Belot:1996:WGR

[425] Gordon Belot. Why General Relativity does need an interpretation. Philosophy of Science, 63:S80–S88, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188514.

Bradie:1996:ORS

[426] Michael Bradie. Ontic realism and scientific explanation. Philosophy of Science, 63:S315–S321, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188541.

Chinn:1996:MMD

[427] Clark A. Chinn and William F. Brewer. Mental models in data interpretation. Philosophy of Science, 63:S211–S219, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188529.

Clark:1996:TTF

[428] Austen Clark. True theories, false colors. Philosophy of Science, 63: S143–S150, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188521.

Clifton:1996:WBW

[429] Rob Clifton. On what being a world takes away. Philosophy of Science, 63: S151–S158, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188522.

Cooper:1996:TMB

[430] Gregory Cooper. Theoretical modeling and biological laws. Philosophy of Science, 63:S28–S35, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188508.

Darden:1996:Pa

[431] Lindley Darden. Preface. Philosophy of Science, 63:v–vi, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188504.

Dedrick:1996:CCR

[432] Don Dedrick. Can colour be reduced to anything? Philosophy of Science, 63:S134–S142, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188520.

DeVito:1996:CIC

[433] Scott DeVito. Completeness and indeterministic causation. Philosophy of Science, 63:S177–S184, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188525.

Dickson:1996:LFM

[434] Michael Dickson. Logical foundations for modal interpretations of quantum mechanics. Philosophy of Science, 63:S322–S329, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188542.

Douven:1996:PE

[435] Igor Douven. A paradox for empiricism (?). Philosophy of Science, 63: S290–S297, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188538.

Gorham:1996:SIR

[436] Geoffrey Gorham. Similarity as an intertheory relation. Philosophy of Science, 63:S220–S229, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188530.

Griffiths:1996:DPS

[437] Paul E. Griffiths. Darwinism, process structuralism, and natural kinds. Philosophy of Science, 63:S1–S9, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/188505.

Guarini:1996:TPS

[438] Marcello Guarini. Tensor products and split-level architecture: Foundational issues in the classicism-connectionism debate. Philosophy of Science, 63:S239–S247, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188532.

Hall:1996:ECV

[439] Richard J. Hall. The evolution of color vision without colors. Philosophy of Science, 63:S125–S133, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188519.

Harper:1996:IPG

[440] William Harper and Robert DiSalle. Inferences from phenomena in gravitational physics. Philosophy of Science, 63:S46–S54, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188510.

Hemmo:1996:PWM

[441] Meir Hemmo. Possible worlds in the modal interpretation. Philosophy of Science, 63:S330–S337, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188543.

Huggett:1996:EMI

[442] Nick Huggett and Robert Weingard. Exposing the machinery of infinite renormalization. Philosophy of Science, 63:S159–S167, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188523.

Kadane:1996:WSB

[443] Joseph B. Kadane, Mark J. Schervish, and Teddy Seidenfeld. When several Bayesians agree that there will be no reasoning to a foregone conclusion. Philosophy of Science, 63:S281–S289, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188537.

Kukla:1996:AES

[444] Andre Kukla. Antirealist explanations of the success of science. Philosophy of Science, 63:S298–S305, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188539.

Liu:1996:GIC

[445] Chuang Liu. Gauge invariance, Cauchy problem, indeterminism, and symmetry breaking. Philosophy of Science, 63:S71–S79, September 1996. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188513.

Melnyk:1996:TRE

[446] Andrew Melnyk. Testament of a recovering eliminativist. Philosophy of Science, 63:S185–S193, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188526.

Mikkelson:1996:SLA

[447] Gregory M. Mikkelson. Stretched lines, averted leaps, and excluded competition: A theory of scientific counterfactuals. Philosophy of Science, 63: S194–S201, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188527.

Millstein:1996:RDO

[448] Roberta L. Millstein. Random drift and the omniscient viewpoint. Philosophy of Science, 63:S10–S18, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188506.

Parascandola:1996:EAE

[449] Mark Parascandola. Evidence and association: Epistemic confusion in toxic tort law. Philosophy of Science, 63:S168–S176, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188524.

Psillos:1996:SRP

[450] Stathis Psillos. Scientific realism and the ’pessimistic induction’. Philosophy of Science, 63:S306–S314, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188540.

Ruetsche:1996:VFP

[451] Laura Ruetsche. Van Fraassen on preparation and measurement. Philosophy of Science, 63:S338–S346, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188544.

Rynasiewicz:1996:TSS

[452] Robert Rynasiewicz. Is there a syntactic solution to the hole problem? Philosophy of Science, 63:S55–S62, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188511.

Sargent:1996:UVP

[453] Pauline Sargent. On the use of visualizations in the practice of science. Philosophy of Science, 63:S230–S238, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188531.

Schmaus:1996:ECM

[454] Warren Schmaus. The empirical character of methodological rules. Philosophy of Science, 63:S98–S106, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188516.

Schwitzgebel:1996:TCR

[455] Eric Schwitzgebel. Theories in children and the rest of us. Philosophy of Science, 63:S202–S210, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188528.

Shaffer:1996:UBS

[456] Nancy E. Shaffer. Understanding bias in scientific practice. Philosophy of Science, 63:S89–S97, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188515.

Siegel:1996:IRN

[457] Harvey Siegel. Instrumental rationality and naturalized philosophy of science. Philosophy of Science, 63:S116–S124, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188518.

Staley:1996:NSH

[458] Kent W. Staley. Novelty, severity, and history in the testing of hypotheses: The case of the top quark. Philosophy of Science, 63:S248–S255, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188533.

Vance:1996:HAG

[459] Russell E. Vance. Heroic antireductionism and genetics: A tale of one science. Philosophy of Science, 63:S36–S45, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188509.

Vineberg:1996:CIM

[460] Susan Vineberg. Confirmation and the indispensability of mathematics to science. Philosophy of Science, 63:S256–S263, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188534.

Weinstein:1996:SCS

[461] Steven Weinstein. Strange couplings and space-time structure. Philosophy of Science, 63:S63–S70, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188512.

Papineau:1996:TDT

[462] David Papineau. Theory-dependent terms. Philosophy of Science, 63(1): 1–20, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188222.

Milne:1996:OTM

[463] Peter Milne. log[P (h/eb)/P (h/b)] is the one true measure of confirmation. Philosophy of Science, 63(1):21–26, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188223.

Clarke:1996:DAI

[464] Murray Clarke. Darwinian algorithms and indexical representation. Philosophy of Science, 63(1):27–48, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188224.

Barrett:1996:EAA

[465] Jeffrey A. Barrett. Empirical adequacy and the availability of reliable records in quantum mechanics. Philosophy of Science, 63(1):49–64, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188225.

Rappaport:1996:IBE

[466] Steven Rappaport. Inference to the best explanation: Is it really different from Mill’s methods? Philosophy of Science, 63(1):65–80, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188226.

Seager:1996:NQE

[467] William Seager. A note on the ‘quantum eraser’. Philosophy of Science, 63(1):81–90, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188227.

Leeds:1996:NVF

[468] Stephen Leeds and Richard Healey. A note on van Fraassen’s modal interpretation of quantum mechanics. Philosophy of Science, 63(1):91–104, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188228.

Butler:1996:CCP

[469] Keith Butler. Content, causal powers, and context. Philosophy of Science, 63(1):105–114, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188229.

VanDerSteen:1996:SNS

[470] Wim J. Van Der Steen. Screening-off and natural selection. Philosophy of Science, 63(1):115–121, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188230.

Richardson:1996:BRB

[471] Robert C. Richardson. Book review: Adaptation and Environment by Robert N. Brandon. Philosophy of Science, 63(1):122–136, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188231.

McLaughlin:1996:BRB

[472] Brian P. McLaughlin. Book review: The Computational Brain by Patricia S. Churchland; Terrence J. Sejnowski. Philosophy of Science, 63(1):137– 139, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188232.

Wilson:1996:BRB

[473] Bradley E. Wilson. Book review: Instrumental Biology, or the Disunity of Science by Alexander Rosenberg. Philosophy of Science, 63(1):139– 141, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188233.

Baird:1996:BRB

[474] Davis Baird. Book review: The Creation of Scientific Effects: Heinrich Hertz and Electric Waves by Jed Z. Buchwald. Philosophy of Science, 63(1):141–143, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188234.

Dupre:1996:BRB

[475] John Dupre. Book review: From a Biological Point of View: Essays in Evolutionary Philosophy by Elliott Sober. Philosophy of Science, 63(1): 143–145, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188235.

Bishop:1996:BRB

[476] Michael Bishop. Book review: Rethinking Objectivity by Allan Megill. Philosophy of Science, 63(1):145–146, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188236.

Gifford:1996:BRB

[477] Fred Gifford. Book review: Discovery and Explanation in Biology and Medicine by Kenneth F. Schaffner. Philosophy of Science, 63(1):147– 148, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188237.

Anonymous:1996:BMa

[478] Anonymous. Back matter. Philosophy of Science, 63(1):??, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188238.

Anonymous:1996:FMa

[479] Anonymous. Front matter. Philosophy of Science, 63(1):??, March 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188221.

Maher:1996:SOC

[480] Patrick Maher. Subjective and objective confirmation. Philosophy of Science, 63(2):149–174, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188467. See [481].

Achinstein:1996:SER

[481] Peter Achinstein. Swimming in evidence: A reply to Maher. Philosophy of Science, 63(2):175–182, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188468. See [480].

Beller:1996:RAC

[482] Mara Beller. The rhetoric of antirealism and the Copenhagen spirit. Philosophy of Science, 63(2):183–204, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188469.

Needham:1996:MOE

[483] Paul Needham. Macroscopic objects: An exercise in Duhemian ontology. Philosophy of Science, 63(2):205–224, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188470.

Lange:1996:LAL

[484] Marc Lange. Life, “artificial life,” and scientific explanation. Philosophy of Science, 63(2):225–244, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188471.

Newman:1996:ESA

[485] David V. Newman. Emergence and strange attractors. Philosophy of Science, 63(2):245–261, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188472.

Mayr:1996:WSW

[486] Ernst Mayr. What is a species, and what is not? Philosophy of Science, 63(2):262–277, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188473.

Shelley:1996:VAR

[487] Cameron Shelley. Visual abductive reasoning in archaeology. Philosophy of Science, 63(2):278–301, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188474.

Huggett:1996:BRB

[488] Nick Huggett and Robert Weingard. Book review: Interpretive Introduction to Quantum Field Theory by Paul Teller. Philosophy of Science, 63(2):302–314, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188475.

Anonymous:1996:BMb

[489] Anonymous. Back matter. Philosophy of Science, 63(2):??, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188476.

Anonymous:1996:FMb

[490] Anonymous. Front matter. Philosophy of Science, 63(2):??, June 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188466.

Brandon:1996:ICE

[491] Robert N. Brandon and Scott Carson. The indeterministic character of evolutionary theory: No “no hidden variables proof” but no Room for determinism either. Philosophy of Science, 63(3):315–337, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188099.

Dwyer:1996:BL

[492] Susan Dwyer and Paul M. Pietroski. Believing in language. Philosophy of Science, 63(3):338–373, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188100.

Karakostas:1996:BSA

[493] Vassilios Karakostas. On the Brussels School’s Arrow of Time in quantum theory. Philosophy of Science, 63(3):374–400, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188101.

Barnes:1996:TMP

[494] Eric Barnes. Thoughts on Maher’s predictivism. Philosophy of Science, 63 (3):401–410, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188102.

Weber:1996:FMP

[495] Marcel Weber. Fitness made physical: The supervenience of biological concepts revisited. Philosophy of Science, 63(3):411–431, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188103.

Fales:1996:PCA

[496] Evan Fales. Plantinga’s case against naturalistic epistemology. Philosophy of Science, 63(3):432–451, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188104.

Rosenberg:1996:BRB

[497] Alex Rosenberg. Book review: Philosophy of Biology and His Philosophy of Biology by Elliot Sober. Philosophy of Science, 63(3):452–464, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188105. See comments

Sober:1996:SCR

[498] Elliott Sober. Some comments on Rosenberg’s review. Philosophy of Science, 63(3):465–469, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188106. See [497].

Callender:1996:TBT

[499] Craig Callender and Robert Weingard. Time, Bohm’s theory, and quantum cosmology. Philosophy of Science, 63(3):470–474, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188107.

Deltete:1996:BRB

[500] Robert J. Deltete. Book review: Reading the Book of Nature by Peter Kosso. Philosophy of Science, 63(3):475–476, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188108.

Horgan:1996:BRB

[501] Terence Horgan. Book review: The Engine of Reason, the Seat of the Soul: A Philosophical Journey into the Brain by Paul M. Churchland. Philosophy of Science, 63(3):476–478, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188109.

Wayne:1996:BRB

[502] Andrew Wayne. Book review: Quantum Mechanics: Historical Contingency and the Copenhagen Hegemony by James T. Cushing. Philosophy of Science, 63(3):478–480, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188110.

Liston:1996:BRB

[503] Michael Liston. Book review: Understanding the Infinite by Shaughan Lavine. Philosophy of Science, 63(3):480–482, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188111.

Slapin:1996:BRB

[504] Steven Slapin. Book review: The Private Science of Louis Pasteur by Gerald L. Geison. Philosophy of Science, 63(3):482–483, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188112.

Anonymous:1996:BMc

[505] Anonymous. Back matter. Philosophy of Science, 63(3):484, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188113.

Anonymous:1996:FMc

[506] Anonymous. Front matter. Philosophy of Science, 63(3):??, September 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188098.

Gopnik:1996:SC

[507] Alison Gopnik. The scientist as child. Philosophy of Science, 63(4):485– 514, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188064. See

commentary [512].

Carey:1996:SCK

[508] Susan Carey and Elizabeth Spelke. Science and core knowledge. Philosophy of Science, 63(4):515–533, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188065.

Fine:1996:SCP

[509] Arthur Fine. Science as Child’s play: Tales from the crib. Philosophy of Science, 63(4):534–537, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188066.

Giere:1996:SA

[510] Ronald N. Giere. The scientist as adult. Philosophy of Science, 63(4): 538–541, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188067.

Nersessian:1996:CP

[511] Nancy J. Nersessian. Child’s play. Philosophy of Science, 63(4):542–546, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188068.

Solomon:1996:CAG

[512] Miriam Solomon. Commentary on Alison Gopnik’s “The Scientist as Child”. Philosophy of Science, 63(4):547–551, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188069. See [507].

Gopnik:1996:RC

[513] Alison Gopnik. Reply to commentators. Philosophy of Science, 63(4): 552–561, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188070.

Sterelny:1996:RG

[514] Kim Sterelny. The return of the group. Philosophy of Science, 63(4): 562–584, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188071.

Dorato:1996:BRN

[515] Mauro Dorato. On becoming, Relativity, and nonseparability. Philosophy of Science, 63(4):585–604, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188072.

Greenwood:1996:FTA

[516] John D. Greenwood. Freud’s ‘tally’ argument, placebo control treatments, and the evaluation of psychotherapy. Philosophy of Science, 63(4):605– 621, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188073. See

replies [517, 518, 605].

Grunbaum:1996:EET

[517] Adolf Grünbaum. Empirical evaluations of theoretical explanations of psychotherapeutic efficacy: A reply to John D. Greenwood. Philosophy of Science, 63(4):622–641, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188074. See [516].

Erwin:1996:EPR

[518] Edward Erwin. The evaluation of psychotherapy: A reply to Greenwood. Philosophy of Science, 63(4):642–651, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188075. See [516].

Fitelson:1996:WHE

[519] Branden Fitelson. Wayne, Horwich, and evidential diversity. Philosophy of Science, 63(4):652–660, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188076.

Myrvold:1996:BDE

[520] Wayne C. Myrvold. Bayesianism and diverse evidence: A reply to Andrew Wayne. Philosophy of Science, 63(4):661–665, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188077. See [362].

Steel:1996:BVD

[521] Daniel Steel. Bayesianism and the value of diverse evidence. Philosophy of Science, 63(4):666–674, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188078.

Anonymous:1996:BMe

[522] Anonymous. Back matter. Philosophy of Science, 63(4):675–676, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188079.

Anonymous:1996:VI

[523] Anonymous. Volume information. Philosophy of Science, 63(4):677–682, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188062.

Anonymous:1996:FMe

[524] Anonymous. Front matter. Philosophy of Science, 63(4):??, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188063.

Anonymous:1996:BMf

[525] Anonymous. Back matter. Philosophy of Science, 64:S480, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188427.

Anonymous:1996:FMf

[526] Anonymous. Front matter. Philosophy of Science, 64:i–xxix, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188383.

Arntzenius:1996:MDT

[527] Frank Arntzenius. Mirrors and the direction of time. Philosophy of Science, 64:S213–S222, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188404.

Beatty:1996:WDB

[528] John Beatty. Why do biologists argue like they do? Philosophy of Science, 64:S432–S443, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188423.

Brandon:1996:DBL

[529] Robert N. Brandon. Does biology have laws? The experimental evidence. Philosophy of Science, 64:S444–S457, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188424.

Burian:1996:CCE

[530] Richard M. Burian. Comments on complexity and experimentation in biology. Philosophy of Science, 64:S279–S291, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188410.

Callender:1996:WPD

[531] Craig Callender. What is ’the problem of the direction of time’? Philosophy of Science, 64:S223–S234, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188405.

Cartwright:1996:MBL

[532] Nancy Cartwright. Models: The blueprints for laws. Philosophy of Science, 64:S292–S303, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188411.

Churchland:1996:TPF

[533] Paul M. Churchland. To Transform the phenomena: Feyerabend, proliferation, and recurrent neural networks. Philosophy of Science, 64:S408– S420, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188421.

Culp:1996:EGP

[534] Sylvia Culp. Establishing genotype/phenotype relationships: Gene targeting as an experimental approach. Philosophy of Science, 64:S268–S278, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188409.

Darden:1996:Pb

[535] Lindley Darden. Preface. Philosophy of Science, 64:vii–viii, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188384.

Franklin:1996:REI

[536] Allan Franklin. Recycling expertise and instrumental loyalty. Philosophy of Science, 64:S42–S52, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188388.

Giere:1996:SIT

[537] Ronald N. Giere. Scientific inference: Two points of view. Philosophy of Science, 64:S180–S184, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188401.

Hands:1996:CEE

[538] D. Wade Hands. Caveat emptor: Economics and contemporary philosophy of science. Philosophy of Science, 64:S107–S116, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188394.

Hausman:1996:CAI

[539] Daniel M. Hausman. Causation, agency, and Independence. Philosophy of Science, 64:S15–S25, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188386.

Hentschel:1996:IIE

[540] Klaus Hentschel. The interplay of instrumentation, experiment, and theory: Patterns emerging from case studies on solar redshift, 1890– 1960. Philosophy of Science, 64:S53–S64, December 1996. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188389.

Howson:1996:EPE

[541] Colin Howson. Error probabilities in error. Philosophy of Science, 64: S185–S194, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188402.

Hughes:1996:MR

[542] R. I. G. Hughes. Models and representation. Philosophy of Science, 64: S325–S336, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188414.

Hull:1996:WWI

[543] David L. Hull. What’s wrong with invisible-hand explanations? Philosophy of Science, 64:S117–S126, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188395.

Humphreys:1996:ES

[544] Paul Humphreys. Emergence, not supervenience. Philosophy of Science, 64:S337–S345, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188415.

Lloyd:1996:FMP

[545] Elisabeth A. Lloyd. Feyerabend, mill, and pluralism. Philosophy of Science, 64:S396–S407, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188420.

Longino:1996:CSS

[546] Helen E. Longino. Comments on science and social responsibility: A role for philosophy of science? Philosophy of Science, 64:S175–S179, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188400.

Mayo:1996:ESL

[547] Deborah G. Mayo. Error statistics and learning from error: Making a virtue of necessity. Philosophy of Science, 64:S195–S212, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188403.

Mirowski:1996:PET

[548] Philip Mirowski. On playing the economics trump card in the philosophy of science: Why it did not work for Michael Polanyi. Philosophy of Science, 64:S127–S138, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188396.

Mitchell:1996:PL

[549] Sandra D. Mitchell. Pragmatic laws. Philosophy of Science, 64:S468–S479, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188426.

Morgan:1996:TAM

[550] Mary S. Morgan. The technology of analogical models: Irving Fisher’s monetary worlds. Philosophy of Science, 64:S304–S314, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188412.

Morrison:1996:PMB

[551] Margaret Morrison. Physical models and biological contexts. Philosophy of Science, 64:S315–S324, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188413.

Orden:1996:FNF

[552] Guy C. Van Orden and Kenneth R. Paap. Functional neuroimages fail to discover pieces of mind in the parts of the brain. Philosophy of Science, 64:S85–S94, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188392.

Preston:1996:FRR

[553] John Preston. Feyerabend’s retreat from realism. Philosophy of Science, 64:S421–S431, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188422.

Price:1996:TSM

[554] Huw Price. Time symmetry in microphysics. Philosophy of Science, 64: S235–S244, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188406.

Rheinberger:1996:ECB

[555] Hans-Jorg Rheinberger. Experimental complexity in biology: Some epistemological and historical remarks. Philosophy of Science, 64:S245–S254, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188407.

Richardson:1996:NAC

[556] Robert C. Richardson. Natural and artificial complexity. Philosophy of Science, 64:S255–S267, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188408.

Rohrlich:1996:CE

[557] Fritz Rohrlich. Cognitive emergence. Philosophy of Science, 64:S346–S358, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188416.

Rosenberg:1996:CPA

[558] Alex Rosenberg. Can physicalist antireductionism compute the embryo? Philosophy of Science, 64:S359–S371, December 1996. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188417.

Sent:1996:EGG

[559] Esther-Mirjam Sent. An Economist’s glance at Goldman’s economics. Philosophy of Science, 64:S139–S148, December 1996. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188397.

Shimony:1996:SHP

[560] Abner Shimony. Some historical and philosophical reflections on science and enlightenment. Philosophy of Science, 64:S1–S14, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188385.

Shrader-Frechette:1996:HFQ

[561] Kristin Shrader-Frechette. Hydrogeology and framing questions having policy consequences. Philosophy of Science, 64:S149–S160, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188398.

Sober:1996:TOL

[562] Elliott Sober. Two outbreaks of lawlessness in recent philosophy of biology. Philosophy of Science, 64:S458–S467, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188425.

Steinle:1996:ENF

[563] Friedrich Steinle. Entering new fields: Exploratory uses of experimentation. Philosophy of Science, 64:S65–S74, December 1996. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188390.

Stufflebeam:1996:PEM

[564] Robert S. Stufflebeam and William Bechtel. PET: Exploring the myth and the method. Philosophy of Science, 64:S95–S106, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188393.

Tiles:1996:SPH

[565] Mary Tiles. Science and the politics of hunger. Philosophy of Science, 64: S161–S174, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188399.

Trumpler:1996:VVP

[566] Maria Trumpler. Verification and variation: Patterns of experimentation in investigations of Galvanism in Germany, 1790–1800. Philosophy of Science, 64:S75–S84, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188391.

vanFraassen:1996:SEF

[567] Bas C. van Fraassen. Sola experientia? — Feyerabend’s refutation of classical empiricism. Philosophy of Science, 64:S385–S395, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188419.

Wimsatt:1996:ARH

[568] William C. Wimsatt. Aggregativity: Reductive heuristics for finding emergence. Philosophy of Science, 64:S372–S384, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188418.

Woodward:1996:EII

[569] Jim Woodward. Explanation, invariance, and intervention. Philosophy of Science, 64:S26–S41, December 1996. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188387.

Humphreys:1997:HPE

[570] Paul Humphreys. How properties Emerge. Philosophy of Science, 64(1): 1–17, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188367.

Healey:1997:NAB

[571] Richard Healey. Nonlocality and the Aharonov–Bohm Effect. Philosophy of Science, 64(1):18–41, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188368. See comment [650].

Sobel:1997:CPW

[572] Jordan Howard Sobel. Cyclical preferences and world Bayesianism. Philosophy of Science, 64(1):42–73, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188369.

Buller:1997:IEP

[573] David J. Buller. Individualism and evolutionary psychology (or: In defense of “narrow” functions). Philosophy of Science, 64(1):74–95, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188370.

LaPorte:1997:EM

[574] Joseph LaPorte. Essential membership. Philosophy of Science, 64(1):96– 112, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188371.

Gemes:1997:ISP

[575] Ken Gemes. Inductive skepticism and the probability calculus i: Popper and Jeffreys on induction and the probability of law-like universal generalizations. Philosophy of Science, 64(1):113–130, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188372.

Shapiro:1997:CV

[576] Lawrence A. Shapiro. A clearer vision. Philosophy of Science, 64(1):131– 153, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188373.

Hardcastle:1997:EEB

[577] Valerie Gray Hardcastle. [explanation] is explanation better. Philosophy of Science, 64(1):154–160, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188374.

Currie:1997:MSM

[578] Gregory Currie and Ian Ravenscroft. Mental simulation and motor imagery. Philosophy of Science, 64(1):161–180, March 1997. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188375.

Solomon:1997:BRB

[579] Miriam Solomon. Book review: Cognition in the Wild by Edwin Hutchins. Philosophy of Science, 64(1):181–182, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/188376.

Koertge:1997:BRB

[580] Noretta Koertge. Book review: The Myth of the Framework: In Defence of Science and Rationality by Karl R. Popper; M. A. Notturno. Philosophy of Science, 64(1):182–184, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188377.

Sklar:1997:BRB

[581] Lawrence Sklar. Book review: In the Wake of Chaos: Unpredictable Order in Dynamic Systems by Stephen H. Kellert. Philosophy of Science, 64(1): 184–185, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188378.

Stephens:1997:BRB

[582] Lynn Stephens. Book review: Rewriting the Soul: Multiple Personality and the Sciences of Memory by Ian Hacking. Philosophy of Science, 64 (1):185–187, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188379.

Hull:1997:BRB

[583] David L. Hull. Book review: A Philosophical Testament by Marjorie Grene. Philosophy of Science, 64(1):187–188, March 1997. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188380.

Grush:1997:BRB

[584] Rick Grush. Book review: The Cognitive Neurosciences by Michael S. Gazzaniga. Philosophy of Science, 64(1):188–190, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188381.

Anonymous:1997:BMa

[585] Anonymous. Back matter. Philosophy of Science, 64(1):??, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188382.

Anonymous:1997:FMa

[586] Anonymous. Front matter. Philosophy of Science, 64(1):??, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188365.

Demopoulos:1997:ORB

[587] William Demopoulos and Philip Kitcher. [obituary: Robert E. Butts]. Philosophy of Science, 64(1):??, March 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188366.

Hellman:1997:BB

[588] Geoffrey Hellman. Bayes and beyond. Philosophy of Science, 64(2):191– 221, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188305. See com-

ments [739].

Mayo:1997:DPB

[589] Deborah G. Mayo. Duhem’s Problem, the Bayesian way, and error statistics, or “what’s belief got to do with it?”. Philosophy of Science, 64(2): 222–244, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188306.

Kelly:1997:LTP

[590] Kevin T. Kelly, Oliver Schulte, and Cory Juhl. Learning theory and the philosophy of science. Philosophy of Science, 64(2):245–267, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188307.

Howson:1997:LI

[591] Colin Howson. A logic of induction. Philosophy of Science, 64(2):268– 290, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188308.

Maher:1997:DDB

[592] Patrick Maher. Depragmatized Dutch book arguments. Philosophy of Science, 64(2):291–305, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188309.

Laudan:1997:HAB

[593] Larry Laudan. How about bust? Factoring explanatory power back into theory evaluation. Philosophy of Science, 64(2):306–316, June 1997. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188310.

Hellman:1997:RMK

[594] Geoffrey Hellman. Responses to Maher, and to Kelly, Schulte, and Juhl. Philosophy of Science, 64(2):317–322, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188311.

Mayo:1997:RHL

[595] Deborah G. Mayo. Response to Howson and Laudan. Philosophy of Science, 64(2):323–333, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188312.

Rescher:1997:HHH

[596] Nicholas Rescher. H_{2}O: Hempel–Helmer–Oppenheim, an episode in the history of scientific philosophy in the 20th Century. Philosophy of Science, 64(2):334–360, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188313.

Buchwald:1997:TK

[597] Jed Z. Buchwald and George E. Smith. Thomas S. Kuhn, 1922–1996. Philosophy of Science, 64(2):361–376, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188314.

Anonymous:1997:BMb

[598] Anonymous. Back matter. Philosophy of Science, 64(2):??, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188315.

Anonymous:1997:FMb

[599] Anonymous. Front matter. Philosophy of Science, 64(2):??, June 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188304.

Kyburg:1997:QMN

[600] Henry E. Kyburg, Jr. Quantities, magnitudes, and numbers. Philosophy of Science, 64(3):377–410, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188317.

Herron:1997:DCF

[601] Timothy Herron, Teddy Seidenfeld, and Larry Wasserman. Divisive conditioning: Further results on dilation. Philosophy of Science, 64(3):411–444, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188318.

Cunningham:1997:TFI

[602] Suzanne Cunningham. Two faces of intentionality. Philosophy of Science, 64(3):445–460, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188319.

Salmon:1997:CER

[603] Wesley C. Salmon. Causality and explanation: A reply to two critiques. Philosophy of Science, 64(3):461–477, September 1997. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188320.

Laudisa:1997:CNA

[604] Federico Laudisa. Contextualism and nonlocality in the algebra of EPR observables. Philosophy of Science, 64(3):478–496, September 1997. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188321.

Greenwood:1997:PCT

[605] John D. Greenwood. Placebo control treatments and the evaluation of psychotherapy: A reply to Grunbaum and Erwin. Philosophy of Science, 64 (3):497–510, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188322.

See [517, 518, 516].

Bandyopadhyay:1997:ICE

[606] Prasanta S. Bandyopadhyay. On an inconsistency in constructive empiricism. Philosophy of Science, 64(3):511–514, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188323.

Hannan:1997:BRB

[607] Barbara Hannan. Book review: Cartesian Psychology and Physical Minds: Individualism and the Sciences of the Mind by Robert A. Wilson. Philosophy of Science, 64(3):515–516, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188324.

Dickson:1997:BRB

[608] Michael Dickson. Book review: Quantum Non-Locality and Relativity: Metaphysical Intimations of Modern Physics by Tim Maudlin. Philosophy of Science, 64(3):516–517, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188325.

Schabas:1997:BRB

[609] Margaret Schabas. Book review: The Values of Precision by M. Norton Wise. Philosophy of Science, 64(3):517–519, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188326.

Daston:1997:BRB

[610] Lorraine Daston. Book review: Discipline and Experience: The Mathematical Way in the Scientific Revolution by Peter Dear. Philosophy of Science, 64(3):519–521, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188327.

Shrader-Frechette:1997:BRB

[611] Kristin Shrader-Frechette. Book review: The Book of Risks: Fascinating Facts about the Chances We Take Every Day by Larry Laudan. Philosophy of Science, 64(3):521–523, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188328.

Anonymous:1997:BMc

[612] Anonymous. Back matter. Philosophy of Science, 64(3):524, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188329.

Anonymous:1997:FMc

[613] Anonymous. Front matter. Philosophy of Science, 64(3):??, September 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188316.

Goldman:1997:SPC

[614] Alvin I. Goldman. Science, publicity, and consciousness. Philosophy of Science, 64(4):525–545, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188559.

Niiniluoto:1997:RIT

[615] Ilkka Niiniluoto. Reference invariance and truthlikeness. Philosophy of Science, 64(4):546–554, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188560.

Charland:1997:RCP

[616] Louis C. Charland. Reconciling cognitive and perceptual theories of emotion: A representational proposal. Philosophy of Science, 64(4):555–579, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188561.

Snyder:1997:DI

[617] Laura J. Snyder. Discoverers’ induction. Philosophy of Science, 64(4): 580–604, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188562.

Glennan:1997:CUS

[618] Stuart S. Glennan. Capacities, universality, and singularity. Philosophy of Science, 64(4):605–626, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188563.

Ramsey:1997:BFP

[619] Jeffry L. Ramsey. Between the fundamental and the phenomenological: The challenge of ‘semi-empirical’ methods. Philosophy of Science, 64(4): 627–653, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188564.

Horvath:1997:SQA

[620] Christopher D. Horvath. Some questions about identifying individuals: Failed intuitions about organisms and species. Philosophy of Science, 64 (4):654–668, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188565.

vanFraassen:1997:MIR

[621] Bas C. van Fraassen. Modal interpretation of repeated measurement: A rejoinder to Leeds and Healey. Philosophy of Science, 64(4):669–676, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188566.

Wagner:1997:OEN

[622] Carl G. Wagner. Old evidence and new explanation. Philosophy of Science, 64(4):677–691, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188567.

Gluck:1997:ICC

[623] Stuart Gluck and Steven Gimbel. An intervening cause counterexample to Railton’s DNP model of explanation. Philosophy of Science, 64(4): 692–697, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188568.

Rottschaefer:1997:AFA

[624] William A. Rottschaefer. Adaptational functional ascriptions in evolutionary biology: A critique of Schaffner’s views. Philosophy of Science, 64 (4):698–713, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188569.

Anonymous:1997:BMd

[625] Anonymous. Back matter. Philosophy of Science, 64(4):714–715, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188570.

Anonymous:1997:VI

[626] Anonymous. Volume information. Philosophy of Science, 64(4):717–722, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188557.

Anonymous:1997:FMd

[627] Anonymous. Front matter. Philosophy of Science, 64(4):??, December 1997. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188558.

Glymour:1998:WWW

[628] Clark Glymour. What went wrong? Reflections on science by observation and the bell curve. Philosophy of Science, 65(1):1–32, March 1998.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188173.

Bishop:1998:FRH

[629] Michael A. Bishop and Stephen P. Stich. The flight to reference, or how not to make progress in the philosophy of science. Philosophy of Science, 65(1):33–49, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188174.

Kronz:1998:NQC

[630] Frederick M. Kronz. Nonseparability and quantum chaos. Philosophy of Science, 65(1):50–75, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188175.

DArms:1998:GTE

[631] Justin D’Arms, Robert Batterman, and Krzyzstof Gorny. Game theoretic explanations and the evolution of justice. Philosophy of Science, 65(1):76– 102, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188176.

Ereshefsky:1998:SPA

[632] Marc Ereshefsky. Species pluralism and Anti-realism. Philosophy of Science, 65(1):103–120, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188177.

Huttemann:1998:LD

[633] Andreas Huttemann. Laws and dispositions. Philosophy of Science, 65 (1):121–135, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188178.

Bozin:1998:ACO

[634] Dragana Bozin. Alternative combining operations in extensive measurement. Philosophy of Science, 65(1):136–150, March 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188179.

Teller:1998:HWR

[635] Paul Teller. On Huggett and Weingard’s review of an Interpretive Introduction to Quantum Field Theory: Continuing the discussion. Philosophy of Science, 65(1):151–161, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188180.

Dowe:1998:BRB

[636] Phil Dowe. Book review: The Facts of Causation by D. H. Mellor. Philosophy of Science, 65(1):162–170, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188181.

Solomon:1998:BRB

[637] Miriam Solomon. Book review: Born to Rebel by Frank Sulloway. Philosophy of Science, 65(1):171–181, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188182.

Anonymous:1998:BMa

[638] Anonymous. Back matter. Philosophy of Science, 65(1):182, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188183.

Anonymous:1998:FMa

[639] Anonymous. Front matter. Philosophy of Science, 65(1):??, March 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188172.

Batterman:1998:WES

[640] Robert W. Batterman. Why equilibrium statistical mechanics works: Universality and the renormalization group. Philosophy of Science, 65(2):183– 208, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188256.

Schaffner:1998:GBD

[641] Kenneth F. Schaffner. Genes, behavior, and developmental emergentism: One process, indivisible? Philosophy of Science, 65(2):209–252, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188257. See commentary [643].

Griffiths:1998:WDC

[642] Paul E. Griffiths and Robin D. Knight. What is the developmentalist challenge? Philosophy of Science, 65(2):253–258, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188258.

Gilbert:1998:WCK

[643] Scott F. Gilbert and Erik M. Jorgensen. Wormwholes: A commentary on K. F. Schaffner’s “Genes, Behavior, and Developmental Emergentism”. Philosophy of Science, 65(2):259–266, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188259. See [641].

Wimsatt:1998:SSP

[644] William C. Wimsatt. Simple systems and phylogenetic diversity. Philosophy of Science, 65(2):267–275, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188260.

Schaffner:1998:MOB

[645] Kenneth F. Schaffner. Model organisms and behavioral genetics: A rejoinder. Philosophy of Science, 65(2):276–288, June 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188261.

Hintikka:1998:RSM

[646] Jaakko Hintikka. Ramsey sentences and the meaning of quantifiers. Philosophy of Science, 65(2):289–305, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188262.

Berger:1998:USW

[647] Ruth Berger. Understanding science: Why causes are not enough. Philosophy of Science, 65(2):306–332, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188263.

Reisch:1998:PLE

[648] George A. Reisch. Pluralism, logical empiricism, and the problem of pseudoscience. Philosophy of Science, 65(2):333–348, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188264.

Halpin:1998:LTH

[649] John F. Halpin. Lewis, Thau, and Hall on chance and the best-system account of law. Philosophy of Science, 65(2):349–360, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188265.

Maudlin:1998:HAB

[650] Tim Maudlin. Healey on the Aharonov–Bohm Effect. Philosophy of Science, 65(2):361–368, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188266. See [571].

Richardson:1998:BRB

[651] Alan Richardson. Book review: Otto Neurath: Philosophy between Science and Politics by Nancy Cartwright; Jordi Cat; Lola Fleck; Thomas E. Uebel. Philosophy of Science, 65(2):369–370, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188267.

Sintonen:1998:BRB

[652] Matti Sintonen. Book review: Error and the Growth of Experimental Knowledge by Deborah Mayo. Philosophy of Science, 65(2):370–372, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188268.

McGowan:1998:BRB

[653] Mary Kate McGowan. Book review: Reading Putnam by Peter Clark; Bob Hale. Philosophy of Science, 65(2):372–373, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188269.

Hardcastle:1998:BRB

[654] Gary L. Hardcastle. Book review: Osiris, Volume 10: Constructing Knowledge in the History of Science by Arnold Thackray. Philosophy of Science, 65(2):373–375, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188270.

Millikan:1998:BRB

[655] Ruth Garrett Millikan. Book review: Complexity and the Function of Mind in Nature by Peter Godfrey-Smith. Philosophy of Science, 65(2): 375–377, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188271.

Kornblith:1998:BRB

[656] Hilary Kornblith. Book review: Explaining Attitudes: A Practical Approach to the Mind by Lynne Rudder Baker. Philosophy of Science, 65(2): 377–379, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188272.

Anonymous:1998:BMb

[657] Anonymous. Back matter. Philosophy of Science, 65(2):380, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188273.

Anonymous:1998:FMb

[658] Anonymous. Front matter. Philosophy of Science, 65(2):??, June 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188255.

Suppe:1998:SSP

[659] Frederick Suppe. The structure of a scientific paper. Philosophy of Science, 65(3):381–405, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 188275. See comment [661].

Lipton:1998:BES

[660] Peter Lipton. The best explanation of a scientific paper. Philosophy of Science, 65(3):406–410, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188276.

Franklin:1998:CBS

[661] Allan Franklin and Colin Howson. Comment on “The Structure of a Scientific Paper” by Frederick Suppe. Philosophy of Science, 65(3):411– 416, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188277. See

[659].

Suppe:1998:RC

[662] Frederick Suppe. Reply to commentators. Philosophy of Science, 65(3): 417–424, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188278.

Hitchcock:1998:CCP

[663] Christopher Hitchcock. The common cause principle in historical linguistics. Philosophy of Science, 65(3):425–447, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188279.

Glymour:1998:CNP

[664] Bruce Glymour. Contrastive, non-probabilistic statistical explanations. Philosophy of Science, 65(3):448–471, September 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188280.

Harms:1998:UIT

[665] William F. Harms. The use of information theory in epistemology. Philosophy of Science, 65(3):472–501, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188281.

Laraudogoitia:1998:SRH

[666] Jon Perez Laraudogoitia. Some relativistic and higher order supertasks. Philosophy of Science, 65(3):502–517, September 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188282.

Thalos:1998:TSA

[667] Mariam Thalos. The Trouble with superselection accounts of measurement. Philosophy of Science, 65(3):518–544, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188283.

Anonymous:1998:BMc

[668] Anonymous. Back matter. Philosophy of Science, 65(3):??, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188284.

Anonymous:1998:FMc

[669] Anonymous. Front matter. Philosophy of Science, 65(3):??, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188274.

Skyrms:1998:SCR

[670] Brian Skyrms. Subjunctive conditionals and revealed preference. Philosophy of Science, 65(4):545–574, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188573.

Joyce:1998:NVP

[671] James M. Joyce. A nonpragmatic vindication of probabilism. Philosophy of Science, 65(4):575–603, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188574.

Egan:1998:MI

[672] Frances Egan. The Moon illusion. Philosophy of Science, 65(4):604–623, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188575.

Koperski:1998:MCC

[673] Jeffrey Koperski. Models, confirmation, and chaos. Philosophy of Science, 65(4):624–648, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188576.

Steel:1998:WWM

[674] Daniel Steel. Warfare and western manufactures: A case study of explanation in anthropology. Philosophy of Science, 65(4):649–671, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188577. See [675, 676].

Jones:1998:UDH

[675] Todd Jones. Unification, deduction, and history: A reply to Steel. Philosophy of Science, 65(4):672–681, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/188578. See [674, 676].

Steel:1998:RJ

[676] Daniel Steel. A reply to Jones. Philosophy of Science, 65(4):682–687, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188579. See [675, 674].

Vranas:1998:EES

[677] Peter B. M. Vranas. Epsilon-ergodicity and the success of equilibrium statistical mechanics. Philosophy of Science, 65(4):688–708, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188580.

Godfrey-Smith:1998:MED

[678] Peter Godfrey-Smith. Maternal effects: On Dennett and Darwin’s dangerous idea. Philosophy of Science, 65(4):709–720, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188581.

Sargent:1998:BRB

[679] Rose-Mary Sargent. Book review: The Mangle of Practice: Time, Agency, and Science by Andrew Pickering. Philosophy of Science, 65(4):721–722, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188582.

Pine:1998:BRB

[680] Ronald C. Pine. Book review: A Social History of Truth: Civility and Science in Seventeenth-Century England by Steven Shapin. Philosophy of Science, 65(4):722–725, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188583.

Mendola:1998:BRB

[681] Joseph Mendola. Book review: For the Sake of the Argument: Ramsey Test Conditionals, Inductive Inference, and Nonmonotonic Reasoning by Isaac Levi. Philosophy of Science, 65(4):725–726, December 1998. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188584.

Carroll:1998:BRB

[682] John W. Carroll. Book review: Decision Theory as Philosophy by Mark Kaplan. Philosophy of Science, 65(4):727–728, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188585.

Owens:1998:BRB

[683] Joseph Owens. Book review: Rule-Following and Realism by Gary Ebbs. Philosophy of Science, 65(4):728–730, December 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188586.

Seta:1998:BRB

[684] Marco Del Seta. Book review: On the Reliability of Economic Models by Daniel Little. Philosophy of Science, 65(4):730–732, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188587.

Anonymous:1998:BMd

[685] Anonymous. Back matter. Philosophy of Science, 65(4):733–734, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188588.

Anonymous:1998:VI

[686] Anonymous. Volume information. Philosophy of Science, 65(4):735–740, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188571.

Anonymous:1998:FMd

[687] Anonymous. Front matter. Philosophy of Science, 65(4):??, December 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188572.

Allchin:1998:DWS

[688] Douglas Allchin. Do we see through a social microscope?: Credibility as a vicarious selector. Philosophy of Science, 66:S287–S298, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188778.

Anonymous:1998:BMe

[689] Anonymous. Back matter. Philosophy of Science, 66:S553–S568, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188798.

Anonymous:1998:FMe

[690] Anonymous. Front matter. Philosophy of Science, 66:i–v, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188755.

Appelbaum:1998:DIC

[691] Irene Appelbaum. The dogma of isomorphism: A case study from speech perception. Philosophy of Science, 66:S250–S259, September 1998. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188775.

Backe:1998:LPR

[692] Andrew Backe. The likelihood principle and the reliability of experiments. Philosophy of Science, 66:S354–S361, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188783.

Bandyopadhyay:1998:CFP

[693] Prasanta S. Bandyopadhyay and Robert J. Boik. The curve fitting problem: A Bayesian rejoinder. Philosophy of Science, 66:S390–S402, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188786.

Bartels:1998:OET

[694] Andreas Bartels. Objects or events?: Towards an ontology for quantum field theory. Philosophy of Science, 66:S170–S184, September 1998. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188769.

Bartha:1998:NOK

[695] Paul Bartha and Christopher Hitchcock. No one knows the date or the hour: An unorthodox application of Rev. Bayes’s Theorem. Philosophy of Science, 66:S339–S353, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188782.

Belousek:1998:BTN

[696] Darrin W. Belousek. Bell’s Theorem, nonseparability, and spacetime individuation in quantum mechanics. Philosophy of Science, 66:S28–S46, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188759.

Bueno:1998:ECQ

[697] Otávio Bueno. Empiricism, conservativeness, and quasi-truth. Philosophy of Science, 66:S474–S485, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188792.

Catton:1998:PDI

[698] Philip Catton. Problems with the deductivist image of scientific reasoning. Philosophy of Science, 66:S452–S473, September 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188791.

Christensen:1998:OKB

[699] Wayne D. Christensen and Clifford A. Hooker. The organization of knowledge: Beyond Campbell’s evolutionary epistemology. Philosophy of Science, 66:S237–S249, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188774.

Clifton:1998:CSR

[700] Rob Clifton and Laura Ruetsche. Changing the subject: Redei on causal dependence and screening off in relativistic quantum field theory. Philosophy of Science, 66:S156–S169, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188768.

Cordero:1998:GTB

[701] Alberto Cordero. Are GRW tails as bad as they say? Philosophy of Science, 66:S59–S71, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188761.

Curiel:1998:ASS

[702] Erik Curiel. The analysis of singular spacetimes. Philosophy of Science, 66: S119–S145, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188766.

Dickson:1998:LET

[703] Michael Dickson. The light at the end of the tunneling: Observation and underdetermination. Philosophy of Science, 66:S47–S58, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188760.

Doring:1998:WBP

[704] Frank Döring. Why Bayesian psychology is incomplete. Philosophy of Science, 66:S379–S389, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188785.

Douven:1998:IBE

[705] Igor Douven. Inference to the best explanation made coherent. Philosophy of Science, 66:S424–S435, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188789.

Dowe:1998:CQT

[706] Phil Dowe. The conserved quantity theory of causation and chance raising. Philosophy of Science, 66:S486–S501, September 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188793.

Fitelson:1998:PBM

[707] Branden Fitelson. The plurality of Bayesian measures of confirmation and the problem of measure sensitivity. Philosophy of Science, 66:S362–S378, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188784.

Freedman:1998:LNA

[708] Karyn Freedman. Laudan’s naturalistic axiology. Philosophy of Science, 66:S526–S537, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188796.

Glymour:1998:PRS

[709] Bruce Glymour. Is pure R-selection really selection? Philosophy of Science, 66:S185–S195, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188770.

Grantham:1998:EPP

[710] Todd A. Grantham. Explanatory pluralism in paleobiology. Philosophy of Science, 66:S223–S236, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188773.

Hintikka:1998:IE

[711] Jaakko Hintikka and Ilpo Halonen. Interpolation as explanation. Philosophy of Science, 66:S414–S423, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188788.

Howard:1998:P

[712] Don Howard. Preface. Philosophy of Science, 66:vii–viii, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188756.

Hyder:1998:HNC

[713] David Jalal Hyder. Helmholtz’s naturalized conception of geometry and his spatial theory of signs. Philosophy of Science, 66:S273–S286, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188777.

Iseda:1998:UNS

[714] Tetsuji Iseda. Use-novelty, severity, and a systematic neglect of relevant alternatives. Philosophy of Science, 66:S403–S413, September 1998. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188787.

Landry:1998:CTL

[715] Elaine Landry. Category theory: The language of mathematics. Philosophy of Science, 66:S14–S27, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188758.

Laudisa:1998:NNC

[716] Federico Laudisa. A note on nonlocality, causation, and Lorentz invariance. Philosophy of Science, 66:S72–S81, September 1998. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188762.

Liu:1998:EEC

[717] Chuang Liu. Explaining the emergence of cooperative phenomena. Philosophy of Science, 66:S92–S106, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188764.

Monton:1998:VFR

[718] Bradley Monton. Van Fraassen and Ruetsche on preparation and measurement. Philosophy of Science, 66:S82–S91, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188763.

Niiniluoto:1998:DA

[719] Ilkka Niiniluoto. Defending abduction. Philosophy of Science, 66:S436– S451, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188790.

Peressini:1998:APM

[720] Anthony Peressini. Applying pure mathematics. Philosophy of Science, 66:S1–S13, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188757.

Risjord:1998:NSA

[721] Mark Risjord. No strings attached: Functional and intentional action explanations. Philosophy of Science, 66:S299–S313, September 1998. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188779.

Roberts:1998:LNI

[722] John Roberts. “Laws of Nature” as an indexical term: A reinterpretation of Lewis’s best-system analysis. Philosophy of Science, 66:S502–S511, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188794.

Schmaus:1998:FMS

[723] Warren Schmaus. Functionalism and the meaning of social facts. Philosophy of Science, 66:S314–S323, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188780.

Schwartz:1998:PFR

[724] Peter H. Schwartz. Proper function and recent selection. Philosophy of Science, 66:S210–S222, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188772.

Skipper:1998:SEE

[725] Robert A. Skipper, Jr. Selection and the extent of explanatory unification. Philosophy of Science, 66:S196–S209, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188771.

Thalos:1998:UD

[726] Mariam Thalos. Units of decision. Philosophy of Science, 66:S324–S338, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188781.

Trout:1998:MRS

[727] J. D. Trout. Measured realism and statistical inference: An explanation for the fast progress of “hard” psychology. Philosophy of Science, 66: S260–S272, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188776.

Uffink:1998:PCC

[728] Jos Uffink. The Principle of the Common Cause faces the Bernstein Paradox. Philosophy of Science, 66:S512–S525, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188795.

vanLith:1998:RCE

[729] Janneke van Lith. Reconsidering the concept of equilibrium in classical statistical mechanics. Philosophy of Science, 66:S107–S118, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188765.

Weinstein:1998:GGT

[730] Steven Weinstein. Gravity and gauge theory. Philosophy of Science, 66: S146–S155, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188767.

Wray:1998:DLS

[731] K. Brad Wray. A defense of Longino’s social epistemology. Philosophy of Science, 66:S538–S552, September 1998. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188797.

Earman:1999:CLI

[732] John Earman and Jesus Mosterin. A critical look at inflationary cosmology. Philosophy of Science, 66(1):1–49, March 1999. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188736.

Maher:1999:ILR

[733] Patrick Maher. Inductive logic and the ravens paradox. Philosophy of Science, 66(1):50–70, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188737.

Weber:1999:ASE

[734] Marcel Weber. The aim and structure of ecological theory. Philosophy of Science, 66(1):71–93, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188738.

Ross:1999:LP

[735] Patricia A. Ross. The limits of physicalism. Philosophy of Science, 66(1): 94–116, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188739.

Slowik:1999:DSR

[736] Edward Slowik. Descartes, spacetime, and relational motion. Philosophy of Science, 66(1):117–139, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188740.

Graves:1999:ISS

[737] Leslie Graves, Barbara L. Horan, and Alex Rosenberg. Is indeterminism the source of the statistical character of evolutionary theory? Philosophy of Science, 66(1):140–157, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188741.

Frisch:1999:VFD

[738] Mathias Frisch. Van Fraassen’s dissolution of Putnam’s model-theoretic argument. Philosophy of Science, 66(1):158–164, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188742.

Kruse:1999:BBC

[739] Michael Kruse. Beyond Bayesianism: Comments on Hellman’s “Bayes and beyond”. Philosophy of Science, 66(1):165–174, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188743. See [588].

Anonymous:1999:BMa

[740] Anonymous. Back matter. Philosophy of Science, 66(1):??, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188744.

Anonymous:1999:FMa

[741] Anonymous. Front matter. Philosophy of Science, 66(1):??, March 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188735.

Bechtel:1999:MRR

[742] William Bechtel and Jennifer Mundale. Multiple realizability revisited: Linking cognitive and neural states. Philosophy of Science, 66(2):175– 207, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188642.

Sarkar:1999:DMP

[743] Sahotra Sarkar and John Stachel. Did Malament prove the nonconventionality of simultaneity in the Special Theory of Relativity? Philosophy of Science, 66(2):208–220, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188643.

Bedard:1999:MOB

[744] Katherine Bedard. Material objects in Bohm’s interpretation. Philosophy of Science, 66(2):221–242, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188644.

Lange:1999:LCS

[745] Marc Lange. Laws, counterfactuals, stability, and degrees of lawhood. Philosophy of Science, 66(2):243–267, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188645.

Shanks:1999:RCC

[746] Niall Shanks and Karl H. Joplin. Redundant complexity: A critical analysis of intelligent design in biochemistry. Philosophy of Science, 66(2): 268–282, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188646.

Wagner:1999:OEN

[747] Carl G. Wagner. Old evidence and new explanation II. Philosophy of Science, 66(2):283–288, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188647.

Hudson:1999:MSN

[748] Robert G. Hudson. Mesosomes: A study in the nature of experimental reasoning. Philosophy of Science, 66(2):289–309, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188648.

Needham:1999:MP

[749] Paul Needham. Macroscopic processes. Philosophy of Science, 66(2):310– 331, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188649.

Anonymous:1999:BMb

[750] Anonymous. Back matter. Philosophy of Science, 66(2):332, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188650.

Anonymous:1999:FMb

[751] Anonymous. Front matter. Philosophy of Science, 66(2):??, June 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188641.

Salmon:1999:SLE

[752] Wesley C. Salmon. The spirit of logical empiricism: Carl G. Hempel’s role in Twentieth-Century philosophy of science. Philosophy of Science, 66(3): 333–350, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188590.

Ismael:1999:SP

[753] Jenann Ismael. Science and the phenomenal. Philosophy of Science, 66 (3):351–369, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188591.

Byrne:1999:CCQ

[754] Alex Byrne and Ned Hall. Chalmers on consciousness and quantum mechanics. Philosophy of Science, 66(3):370–390, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188592.

Tennant:1999:SEF

[755] Neil Tennant. Sex and the evolution of fair-dealing. Philosophy of Science, 66(3):391–414, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188593.

Hardcastle:1999:SPV

[756] Valerie Gray Hardcastle. Scientific papers have various structures. Philosophy of Science, 66(3):415–439, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188594.

Healey:1999:QAR

[757] Richard Healey. Quantum analogies: A reply to Maudlin. Philosophy of Science, 66(3):440–447, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188595.

Morvan:1999:CCC

[758] Pierre Le Morvan. The converse consequence condition and Hempelian qualitative confirmation. Philosophy of Science, 66(3):448–454, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188596.

Glymour:1999:MTT

[759] Clark Glymour. A mind is a terrible thing to waste — critical notice: Jaegwon Kim, Mind in a Physical World. Philosophy of Science, 66(3): 455–471, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188597.

Fitelson:1999:HDD

[760] Branden Fitelson, Christopher Stephens, and Elliott Sober. How not to detect design — critical notice: William A. Dembski, The Design Inference. Philosophy of Science, 66(3):472–488, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188598.

Dupre:1999:BRBa

[761] John Dupré. Book review: How the Mind Works by Steven Pinker. Philosophy of Science, 66(3):489–493, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188599.

Baker:1999:BRB

[762] Lynne Rudder Baker. Book review: Having Thought: Essays in the Metaphysics of Mind by John Haugeland. Philosophy of Science, 66(3):494–495, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188600.

Dickson:1999:BRB

[763] Michael Dickson. Book review: Interpreting the Quantum World by Jeffrey Bub. Philosophy of Science, 66(3):495–496, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188601.

Hall:1999:BRB

[764] Ned Hall. Book review: Causality and Explanation by Wesley C. Salmon. Philosophy of Science, 66(3):497–498, September 1999. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188602.

Scott-Kakures:1999:BRB

[765] Dion Scott-Kakures. Book review: Brainchildren: Essays on Designing Minds by Daniel C. Dennett. Philosophy of Science, 66(3):498–500, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188603.

McLarty:1999:BRB

[766] Colin McLarty. Book review: Real Numbers, Generalizations of the Reals, & Theories of Continua by Philip Ehrlich. Philosophy of Science, 66(3): 500–501, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188604.

Liston:1999:BRB

[767] Michael Liston. Book review: What Is Mathematics, Really? by Reuben Hersh. Philosophy of Science, 66(3):501–502, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188605.

Balaguer:1999:BRB

[768] Mark Balaguer. Book review: Naturalism in Mathematics by Penelope Maddy. Philosophy of Science, 66(3):502–504, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188606.

Dupre:1999:BRBb

[769] John Dupré. Book review: This Is Biology: The Science of the Living World by Ernst Mayr. Philosophy of Science, 66(3):504–506, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188607.

Clarke:1999:BRB

[770] Steve Clarke. Book review: The Disunity of Science: Boundaries Contexts, and Power by Peter Galison; David J. Stump. Philosophy of Science, 66(3):506–507, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188608.

Anonymous:1999:BMc

[771] Anonymous. Back matter. Philosophy of Science, 66(3):508, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188609.

Anonymous:1999:FMc

[772] Anonymous. Front matter. Philosophy of Science, 66(3):??, September 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188589.

Levi:1999:VCV

[773] Isaac Levi. Value commitments, value conflict, and the separability of belief and value. Philosophy of Science, 66(4):509–533, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188747.

Bishop:1999:WTE

[774] Michael A. Bishop. Why thought experiments are not arguments. Philosophy of Science, 66(4):534–541, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188748.

Sober:1999:MRA

[775] Elliott Sober. The multiple realizability argument against reductionism. Philosophy of Science, 66(4):542–564, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188749.

Downes:1999:CSD

[776] Stephen M. Downes. Can scientific development and Children’s cognitive development be the same process? Philosophy of Science, 66(4):565–578, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188750.

Shelley:1999:MAA

[777] Cameron Shelley. Multiple analogies in archaeology. Philosophy of Science, 66(4):579–605, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188751.

Leeds:1999:GAB

[778] Stephen Leeds. Gauges: Aharonov, Bohm, Yang, Healey. Philosophy of Science, 66(4):606–627, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188752.

Walsh:1999:AI

[779] D. M. Walsh. Alternative individualism. Philosophy of Science, 66(4): 628–648, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188753.

Anonymous:1999:BMd

[780] Anonymous. Back matter. Philosophy of Science, 66(4):649–656, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188754.

Anonymous:1999:VI

[781] Anonymous. Volume information. Philosophy of Science, 66(4):651–655, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188745.

Anonymous:1999:FMd

[782] Anonymous. Front matter. Philosophy of Science, 66(4):??, December 1999. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188746.